### 03 Dec 2023 v2

# TransNAS-TSAD: Harnessing Transformers for Multi-Objective Neural Architecture Search in Time Series Anomaly Detection [^1] This document is the results of the research project funded by the National Science Foundation.22footnotemark: 2

Abstract
The surge in real-time data collection across various industries has underscored the need for advanced anomaly detection in both univariate and multivariate time series data. Traditional methods, while comprehensive, often struggle to capture the complex interdependencies in such data. This paper introduces TransNAS-TSAD, a novel framework that synergizes transformer architecture with neural architecture search (NAS), enhanced through NSGA-II algorithm optimization. This innovative approach effectively tackles the complexities of both univariate and multivariate time series, balancing computational efficiency with detection accuracy. Our evaluation reveals that TransNAS-TSAD surpasses conventional anomaly detection models, demonstrating marked improvements in diverse data scenarios. We also propose the Efficiency-Accuracy-Complexity Score (EACS) as a new metric for assessing model performance, emphasizing the crucial balance between accuracy and computational resources. TransNAS-TSAD sets a new benchmark in time series anomaly detection, offering a versatile, efficient solution for complex real-world applications. This research paves the way for future developments in the field, highlighting its potential in a wide range of industry applications.
keywords:
Time Series Anomaly Detection, Transformers in Machine Learning, Neural Architecture Search (NAS), NSGA-II Optimization, Deep Learning for Time Series Analysis, Adversarial Learning Paradigms
[…]
[1]organization=Department of Computer Science, addressline=University of Vermont,postcode=05405,city=Burlington, VT, country=U.S.A
1 Introduction
The prevalence of time series data in various sectors—from finance to healthcare, from infrastructure to manufacturing—highlights its critical role in modern analytics. Such data are often pivotal in uncovering underlying patterns and dependencies that are vital for strategic insights and anomaly detection, which can indicate significant shifts in system behaviors or emerging issues (Brophy et al., (2023); Li et al., (2023)).
Traditional statistical methods have long been the mainstay for anomaly detection in time series data, trusted for their robust mathematical frameworks (Thudumu et al., (2020); Wang et al., (2011); Leadbetter, (1991)). Yet, the advent of Big Data—with its hallmark features of volume, velocity, and variety—has stretched these methods to their limits, revealing shortcomings through high false positive rates and missed detections (Ashabi et al., (2020); Al-Sai et al., (2019)).
In contrast, deep learning has ushered in a paradigm shift, offering enhanced adaptability and performance in anomaly detection (Ismail Fawaz et al., (2019); Ma et al., (2021); Haq et al., (2023)). The transformer architecture, in particular, has emerged as a groundbreaking development due to its self-attention mechanism, enabling significant strides not only in natural language processing but also in time series analysis (Vaswani et al., (2017); Tuli et al., (2022); Kim et al., (2023)).
However, the specific application of transformer models to time series anomaly detection necessitates further exploration to fine-tune their structures and parameters for optimal performance (Arslan et al., (2023)). Our initiative, TransNAS-TSAD, represents a significant shift from traditional transformer models, such as TranAD, to a more dynamic and adaptable framework. This new model integrates the transformer’s architectural strengths with systematic optimization through neural architecture search (NAS), emphasizing the multi-objective optimization strengths of the NSGA-II algorithm (Elsken et al., (2019); Deb et al., (2002)).
TransNAS-TSAD marks a significant evolution from TranAD’s static architecture, primarily through the integration of NSGA-II. This shift, grounded in (Lu et al., (2023)) study, leverages NSGA-II’s proven superiority in exploring complex search spaces more efficiently than recent algorithms. By adeptly balancing performance with computational efficiency, NSGA-II enhances our neural architecture design for anomaly detection, offering improved adaptability and efficiency. (Lu et al., (2020).
1.1 Research Context and Challenges
Adapting anomaly detection models for time series data is a complex task, especially when dealing with multivariate series that present high-dimensional challenges and necessitate sophisticated model architectures (Wei, (2018); Liu et al., (2018)). The selective attention capabilities of the transformer model present a compelling solution, but only if the architectures are tailored to the specific characteristics of the dataset (Lian et al., (2019)).
The marriage of NAS with multi-objective optimization algorithms like NSGA-II represents a strategic evolution in automated model design, offering a methodical path to discovering optimal architectures that consider multiple, often conflicting, objectives (Liu et al., (2021); Xue et al., (2023)).
1.2 Theoretical Foundations and Methodological Approach
Our methodology builds on the interplay between the theoretical principles of evolutionary algorithms and the empirical strengths of deep learning architectures. We aim to craft a robust NAS framework that can intelligently navigate and select transformer architectures with hyperparameters fine-tuned for anomaly detection in time series data (Wang et al., (2021); Ma et al., (2021)). The use of NSGA-II is central to this approach, providing a mechanism to balance the trade-offs between detection accuracy and model complexity (Wang et al., (2021)).
1.3 Guiding Philosophies and Contributions
Our research is guided by the following key principles:
  • 1.
    Adaptation to Time Series Data: We assert the need for specialized models that respect the unique temporal and dimensional characteristics of time series data.
  • 2.
    Maximizing Transformer Utility: We aim to exploit the full capabilities of the transformer architecture, particularly for temporal data interpretation.
  • 3.
    Strategic Optimization via NAS: We employ NAS, especially NSGA-II, as a systematic tool for refining architectures towards optimal model configurations.
Our contributions are as follows:
  1. 1.
    Novel Application of NAS: We leverage NAS to uniquely optimize transformer architectures for the domain of time series anomaly detection.
  2. 2.
    Architectural Parameter Exploration: We conduct extensive experimentation to identify transformer configurations that redefine design norms and improve detection performance.
  3. 3.
    Benchmarking Performance: We set new standards for performance, outperforming existing methods in a variety of datasets.
  4. 4.
    Model Design Efficiency: Our approach yields models that balance computational efficiency with analytical accuracy.
  5. 5.
    Architectural and Search Strategy Synergy: We seamlessly blend the structural advantages of transformers with the adaptive capabilities of NAS, resulting in models adept for time series anomaly detection.
2 Related Work
2.1 Anomaly Detection in Time Series Data
Anomaly detection in time series data has been a focal point of research due to its critical applications across various domains such as finance, healthcare, and industrial automation. Traditional approaches have primarily relied on statistical methods, which assume specific data distributions or apply distance-based metrics (Leadbetter, (1991); Wang et al., (2011)). However, these methods often struggle with high-dimensional, complex time series data inherent in modern applications (Hodge & Austin, (2004); Patcha & Park, (2007)).
With the advent of Big Data, researchers have sought more robust solutions capable of handling large volumes of high-velocity, high-variety data. Machine learning, and specifically deep learning, has emerged as a powerful approach to detect anomalies by learning representations of normal behavior and identifying deviations (Lee et al., (2022); Landauer et al., (2023); Al-amri et al., (2021)).
2.2 Transformers in Time Series Analysis
, The transformer architecture, introduced by Vaswani et al., (2017) has become a cornerstone for sequence modeling tasks, initially within the realm of natural language, processing and subsequently in time series analysis (Li et al., (2019); Tang & Matteson, (2021). Its self-attention mechanism allows it to capture temporal dependencies without the constraints of sequential data processing, distinguishing it from its predecessors like RNNs and LSTMs (Karita et al., (2019); Reza et al., (2022); Katrompas et al., (2022)).
In the domain of time series anomaly detection, transformers have shown promise due to their ability to model complex patterns over time (Tuli et al., (2022); Liu et al., (2021); Wang et al., (2022)). However, the research is still nascent, with considerable scope for optimization and adaptation to various time series characteristics (Luo et al., (2021); Wen et al., (2022)).
2.3 Neural Architecture Search (NAS)
NAS has emerged as a paradigm to automate the design of neural network architectures. Early NAS approaches were computationally intensive, employing reinforcement learning or evolutionary algorithms to explore the design space (Liu et al., (2021); Chen et al., (2019); Chu et al., (2020)). The development of efficient NAS methods has been a recent trend, focused on reducing the computational resources required for the search process (Jin et al., (2019); He et al., (2020)).
2.4 Multi-Objective Optimization in NAS
Multi-objective optimization in NAS aims to find a trade-off between competing objectives, such as model performance and computational cost. NSGA-II, an evolutionary algorithm, has been widely used for this purpose due to its efficiency in finding a diverse set of Pareto optimal solutions (Wang et al., (2020); Lu et al., (2019); Ying et al., (2020)). The application of multi-objective optimization in NAS for time series anomaly detection is still an under-explored area, offering a fertile ground for research (Borchert et al., (2022); Lu et al., (2021)).
2.5 Convergence of NAS and Transformers
The intersection of NAS and transformer models represents the cutting edge of research in the quest for optimal architectures for specific tasks. In time series anomaly detection, this convergence promises architectures that are both performant and efficient, addressing the unique challenges of temporal data (Wen et al., (2022)).
This work builds upon these foundations, aiming to address the identified gaps by leveraging the strengths of NAS and the transformer model to push the boundaries of anomaly detection in time series data.
3 Methodology
In this section, we delineate the comprehensive methodology employed in our study, designed to effectively detect and analyze anomalies in multivariate time series data using the TransNAS-TSAD framework. We commence with a clear definition of the problem, setting the stage for the subsequent methodological steps. This is followed by an in-depth discussion on data refinement techniques, ensuring that the data is appropriately pre-processed for our analysis. Next, we delve into the intricacies of our transformer architecture, highlighting its design and capabilities. A critical component of our approach is the neural architecture search (NAS) process, employing the NSGA-II algorithm for optimizing the transformer model, which we expound upon in detail. This is supplemented by an overview of the evolutionary process that iteratively refines our model architecture. We then transition to discussing advanced anomaly detection techniques, integrating adversarial elements into our model. The section culminates with a description of how we harness the full potential of TransNAS-TSAD for practical anomaly detection applications, illustrating the real-world implications of our research. Through this structured approach, we aim to provide a clear and logical flow of our methodology, ensuring that each step contributes coherently to achieving our research objectives.
3.1 Problem Definition
Our research pioneers the application of a multi-objective approach to anomaly detection in multivariate time series data, utilizing transformer models. The problem is defined as:
Given a sequence of multivariate time series data $S=\{x_{1},x_{2},\ldots,x_{T}\}$, where each point $x_{t}$ is timestamped $t$ and resides in an $m$-dimensional space ($x_{t}\in\mathbb{R}^{m}$), our aims are:
  1. 1.
    Anomaly Detection in Time Series: Using a training time series $S$ and a test series $S^{\prime}$ of length $T^{\prime}$ with similar modalities, we seek to detect anomalies by predicting a binary sequence $Y=\{y_{1},y_{2},\ldots,y_{T^{\prime}}\}$, where $y_{t}$ identifies anomalies at timestamp $t$ in $S^{\prime}$ (1 indicates an anomaly).
  2. 2.
    Anomaly Component Analysis: Our goal extends to identifying anomalous components within each data point of the series, generating a detailed prediction sequence $Y=\{y_{1},y_{2},\ldots,y_{T^{\prime}}\}$, where each $y_{t}$ pinpoints the specific anomalous dimensions at timestamp $t$.
3.1.1 Multi-Objective Transformer Model Optimization
At the heart of our methodology is the use of NSGA-II based neural architecture search (NAS) for optimizing the transformer architecture, focusing on:
  • 1.
    Maximizing Anomaly Detection Accuracy: We aim to enhance the F1 score for precise anomaly detection.
  • 2.
    Minimizing Architectural Complexity: Concurrently, reducing the model’s size and computational demands for efficiency in practical deployment.
This dual-focused approach is a significant advancement in applying transformer models for anomaly detection.
3.1.2 Data Processing and Model Training
Prior to modeling, the time series data undergoes a preprocessing phase where normalization is applied to ensure uniformity of scale across all series components. The normalization process is formulated as follows:
$x_{t}\leftarrow\frac{x_{t}-\min(S)}{\max(S)-\min(S)+\epsilon}$
where $\min(S)$ and $\max(S)$ denote the component-wise minimum and maximum values observed in the training time series $S$, respectively. The term $\epsilon$ is a small constant introduced to prevent division by zero, ensuring numerical stability.
Subsequently, the time series is transformed into a set of overlapping windows, which serve as the input to our model. A context window of size $K$ is defined around each data point $x_{t}$, resulting in:
$W_{t}=\{x_{t-K+1},\dots,x_{t}\}$
where boundary cases are handled by replicating the first available data point to fill the window, maintaining a fixed size of $K$. This approach encapsulates the temporal dependencies inherent in time series data.
3.1.3 Anomaly Scoring and Thresholding
Our model is designed to assign an anomaly score to each context window $W_{t}$, as opposed to direct binary labels. This score is derived from the reconstruction error of the window, where $O_{t}$ denotes the reconstructed output corresponding to $W_{t}$. A dynamic threshold $D$ is established based on the distribution of the scores computed on the training data, facilitating the distinction between normal and anomalous windows. The precise mechanism for determining $D$ will be further elaborated upon in the subsequent sections, focusing on its adaptive nature and how it relates to the detection accuracy.
3.2 TransNAS-TSAD Transformer Architecture
In the TransNAS-TSAD framework, the transformer architecture is specially adapted for time series anomaly detection. This adaptation involves a unique configuration of the standard transformer model to align with the specific characteristics of time series data.
Data Augmentation Mechanisms Within the Model: TransNAS-TSAD integrates specific data augmentation techniques within the model architecture. These techniques, which can be fine-tuned during training, include:
  • 1.
    Gaussian Noise Augmentation (Optional): Adds Gaussian noise to the input sequence for variability and robustness.
  • 2.
    Time Warping & Time Masking Augmentation (Optional): Applies time warping or masking, enhancing the model’s generalization capabilities.
Encoder Operations: The encoder in TransNAS-TSAD is designed with the following operations:
  • 1.
    Linear Embedding (Optional): Transforms the input via linear mapping to a predefined dimensional space.
  • 2.
    Layer Normalization: Normalizes the sequence for stability in subsequent operations.
  • 3.
    Positional Encoding: Uses either sinusoidal or Fourier positional encoding, depending on the architecture optimized through NAS.
  • 4.
    Multi-Head Self Attention: A key feature in TransNAS-TSAD is the multi-head attention mechanism, which is aligned with the number of features in the data, inspired by the TranAD model. It employs the standard scaled-dot product attention mechanism:
    $\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{\text{model}}}}\right)V$ (1)
    The multi-head attention mechanism is formulated as:
    $\text{MultiHeadAtt}(Q,K,V)=\text{Concat}(H_{1},H_{2},\ldots,H_{h})$ (2)
    where $H_{i}=\text{Attention}(Q_{i},K_{i},V_{i})$.
  • 5.
    Feedforward Neural Networks (FFNs) Post-Attention: Following the attention mechanism, the sequence is processed through a series of feedforward neural networks (FFNs). Each FFN layer, consisting of a linear transformation followed by an activation function, introduces non-linearities crucial for complex pattern recognition. The configuration of these FFN layers, particularly the number of layers and the number of neurons within each layer, is dynamically determined within a predefined range in our NAS search space. This optimization, conducted using Optuna, is designed to balance anomaly detection performance with computational efficiency, ensuring that the FFN dimensions are optimally tuned for the specific demands of time series anomaly detection.
Decoder Operations: The decoder mirrors the encoder’s complexity and includes additional transformations that integrate encoder memory, apply multi-head self-attention, and utilize feedforward neural networks, culminating in a final layer for accurate reconstruction. TransNAS-TSAD can utilize either a dual sequential decoder setup or an iterative approach, depending on the optimized configuration for maximum anomaly detection accuracy.
Figure 1: Workflow of the TransNAS-TSAD Process for Time Series Anomaly Detection using Multi-Objective Neural Architecture Search (NAS) with NSGA-II Optimization.
3.3 Neural Architecture Search (NAS) Framework for TransNAS-TSAD
The NAS framework shown in Figure 1 is pivotal in achieving our primary objective of effectively detecting anomalies in multivariate time series data. By employing the NSGA-II algorithm, our NAS process systematically navigates the vast architectural landscape of transformer models. This step is essential in identifying an architecture that not only excels in anomaly detection accuracy but also aligns with the computational constraints of practical deployment scenarios. The rigorous optimization and selection of transformer architectures through NAS directly contribute to enhancing the performance and efficiency of the TransNAS-TSAD framework, thereby fulfilling our goal of developing a robust and adaptable anomaly detection system.
3.3.1 Search Space Definition
The search space outlined in Table 1 within TransNAS-TSAD represents a diverse and extensive array of potential model architectures and hyperparameters. This assortment is meticulously designed to tailor neural network configurations specifically for time-series anomaly detection:
  • 1.
    Architectural Parameters: These parameters encompass the number and types of layers in the encoder and decoder, optimized attention mechanisms for temporal data analysis, dimensions of feedforward networks, and a variety of positional encoding and normalization methods. Collectively, they enable the model to adapt its architecture for different characteristics of time-series data.
  • 2.
    Training Hyperparameters: This category includes a broad spectrum of hyperparameters, such as learning rates, batch sizes that are adaptable to varying computational resources, and dropout rates to ensure effective regularization. Additionally, the window size parameter, ranging from 10 to 30, is crucial in defining the length of the input sequence window, which directly influences how the model processes and learns from temporal patterns. Data augmentation techniques like time warping and masking are also part of the search space, enabling the model to simulate and learn from varied anomaly scenarios effectively.
The comprehensive nature of the search space in TransNAS-TSAD allows for the exploration and optimization of a wide range of models, ensuring that the most suitable architecture and hyperparameter settings are identified for effective anomaly detection in different time-series datasets.
Parameter Type Parameter Name Search Space Description
Training
Learning Rate
$1\times 10^{-5}$ to $1\times 10^{-1}$ (log scale)
Rate at which the model learns.
Dropout Rate
0.1 to 0.5
Regularization to prevent overfitting.
Batch Size
16 to 128 (step of 16)
Number of samples per training step.
Gaussian Noise
$1\times 10^{-4}$ to $1\times 10^{-1}$ (log scale)
Noise added for robustness.
Time Warping
True, False
Augmentation technique for time series.
Time Masking
True, False
Augmentation technique to mask intervals.
Window Size
10 to 30
Length of the input sequence window for the model.
Architectural
Positional Encoding Type
Sinusoidal, Fourier
Encoding type for sequence position.
Dimension Feedforward
8 to 128 (log scale)
Size of the feedforward network.
Encoder Layers
1 to 3
Number of layers in the encoder.
Decoder Layers
1 to 3
Number of layers in the decoder.
Activation Function
ReLU, Leaky ReLU, Sigmoid, Tanh
Non-linearity after each layer.
Attention Type
Scaled Dot Product
Type of attention mechanism.
Number of Attention Heads
Equal to feature dimension
Parallel attention layers.
Use Linear Embedding
True, False
Option to use a linear embedding layer.
Layer Normalization
Layer, Batch, Instance
Type of normalization used.
Self Conditioning
True, False
Conditioning strategy for the model.
Number of FFN Layers
1 to 3
Layers in the feedforward network.
Phase Type
1phase, 2phase, Iterative
Model’s reconstruction and refinement strategy.
Table 1: Neural Architecture Search Space for TransNAS-TSAD
3.3.2 TransNAS-TSAD Evaluation Strategy
Each trial within the TransNAS-TSAD NAS process entails a rigorous evaluation of models generated with varying architectural and training hyperparameters. This evaluation adopts a multi-objective approach, primarily focusing on two critical aspects: the F1 score and the number of model parameters. The F1 score serves as a key indicator of the model’s accuracy in anomaly detection, balancing precision and recall, while the number of parameters gauges the model’s architectural complexity and computational efficiency. The objective is to identify models that not only demonstrate high proficiency in accurately detecting anomalies (as reflected in a high F1 score) but also maintain a streamlined architecture (evidenced by a lower count of parameters). This dual-criteria assessment ensures the selection of models that are both effective in performance and practical in deployment, aligning with the overarching goal of achieving optimal anomaly detection with computational resourcefulness.
3.3.3 Multi-Objective Optimization with NSGA-II
TransNAS-TSAD employs the NSGA-II algorithm to balance the objectives of performance and computational efficiency. The optimization process is defined by the following equations:
$\mathbf{F}(\mathbf{x})=[f_{1}(\mathbf{x}),f_{2}(\mathbf{x})]$ (3)
where $f_{1}(\mathbf{x})$ is the F1 score and $f_{2}(\mathbf{x})$ is a computational resource metric.
Non-dominated sorting and the calculation of crowding distance are integral parts of NSGA-II:
$\text{NonDomSort}(\mathbf{X})=\bigcup_{i=1}^{k}F_{i}$ (4)
$d(\mathbf{x})=\sum_{i=1}^{k}\left(f_{i}(\mathbf{x}^{+})-f_{i}(\mathbf{x}^{-})\right)$ (5)
3.3.4 Pareto Front Exploration
The Pareto front represents the set of non-dominated solutions, providing an optimal trade-off between conflicting objectives:
$\mathcal{P}=\{\mathbf{x}\in\mathbf{X}\mid\nexists\mathbf{y}\in\mathbf{X}:\mathbf{F}(\mathbf{y})\prec\mathbf{F}(\mathbf{x})\}$ (6)
where $\prec$ indicates that $\mathbf{y}$ dominates $\mathbf{x}$.
3.3.5 Utilization of the Pareto Front
The Pareto front derived from the NAS process provides a practical framework for model selection:
  • 1.
    Interpretation and Analysis: The front is analyzed to discern architectural trade-offs, allowing stakeholders to identify models with the desired balance between accuracy and efficiency.
  • 2.
    Informed Selection: Decision-makers can select models that align with specific performance expectations and operational constraints by examining the Pareto front.
  • 3.
    Resource Allocation: Models on the Pareto front inform resource allocation, directing investments toward architectures with the most favorable trade-offs.
Algorithm 1 provides a detailed procedure that encapsulates our approach within the Optuna framework using NSGA-II (Deb et al., (2002)) for the multi-objective optimization. This algorithm elucidates the sequential steps taken from initialization to final selection, highlighting the integration of the evolutionary search process with transformer-specific architecture and hyperparameter tuning.
Input: Training dataset, validation dataset, predefined search space for transformer architectures and hyperparameters
Output: Optimized transformer architecture and hyperparameters for time series anomaly detection
1 Initialization: Set up the Optuna framework with the NSGA-II sampler. Initialize the search with a population of transformer architectures from the predefined search space. Define evolutionary process parameters: population size, number of generations, mutation and crossover rates. Specify evaluation metrics: F1 score, computational resource utilization (model size, training time), precision, recall, and inference time. for each study in Optuna do
2       for each trial in the study do
3             Define the transformer model architectural parameters and training hyperparameters within the search space constraints. Construct and train the transformer model on the Training dataset. Evaluate the model on the Validation dataset using the specified metrics. Update the study with the trial’s objectives, reflecting the model’s performance and computational efficiency. if early stopping criteria are met based on validation performance or predefined resource constraints then
4                   Terminate the trial prematurely.
5             end if
6
7       end for
8      Selection: Use Optuna’s NSGA-II sampler to select architectures for the next generation, focusing on the Pareto-optimal solutions. Crossover and Mutation: Apply Optuna’s genetic operators to evolve the study’s population, exploring new architectural variants.
9 end for
Final Selection: At the end of the NAS process, use Optuna to extract the best transformer architecture and hyperparameters from the Pareto front, considering both performance metrics and computational efficiency.
Algorithm 1 Multi-Objective NAS for Time Series Anomaly Detection with TransNAS-TSAD
3.3.6 Evolutionary Process in TransNAS-TSAD
The evolutionary process within TransNAS-TSAD serves a critical role in aligning our methodology with the overarching goal of refining anomaly detection capabilities. Through iterative refinement and selection of model architectures, this process ensures that our models are not only accurate in identifying anomalies but also evolve to become more efficient and adaptable to various data characteristics. This continuous improvement and adaptation are key to achieving high-performance anomaly detection, which stands at the core of our research objectives.
  • 1.
    Generation Loop: New models are proposed, evaluated, and bred in each generation, with superior architectures refining the Pareto front.
  • 2.
    Early Stopping: The process incorporates early stopping for models that do not show promise, thus conserving computational resources.
  • 3.
    Final Architecture Selection: The optimal architecture is selected from the Pareto front, ensuring an effective balance between the objectives.
3.4 Offline Optimization of Transformer Architecture for Anomaly Detection
Central to TransNAS-TSAD is the offline, strategic optimization of the transformer architecture, a process critical for achieving an effective balance between anomaly detection accuracy and computational efficiency. This optimization, guided by neural architecture search (NAS) using the NSGA-II algorithm, involves:
  1. 1.
    Comprehensive Trial-Based Exploration: In line with Optuna’s best practices, the process involves conducting over 100 trials to explore a wide array of transformer architectures, ensuring a thorough examination of the solution space.
  2. 2.
    Achieving Pareto Front Efficiency: The key focus of NAS is to identify and refine a collection of architectures that constitute the Pareto front, representing a spectrum of models from high-performance, resource-intensive ones to balanced and resource-efficient alternatives.
  3. 3.
    Iterative Evolutionary Process: Through evolutionary algorithms, these architectures undergo continuous adaptation and improvement, exploring innovative configurations and enhancing performance and efficiency with each generation.
3.4.1 Tailoring Architectures to Specific Deployment Needs
Post the offline NAS process, we are equipped with a diverse array of transformer architectures, each optimized for different operational contexts:
  • 1.
    For Resource-Rich Environments: High F1 score models, though computationally demanding, are suitable for scenarios where resource availability is not a constraint.
  • 2.
    In Resource-Limited Settings: The NAS process also yields architectures that are either balanced in terms of performance and resource usage or are specifically optimized for low-resource environments.
This flexibility allows for the deployment of TransNAS-TSAD in various settings, ensuring efficient and effective anomaly detection tailored to the constraints and requirements of different deployment environments.
3.5 Advanced Anomaly Detection with Adversarial Elements in TransNAS-TSAD
TransNAS-TSAD represents a significant advancement in the field of anomaly detection in time series data, strategically incorporating and extending adversarial learning paradigms. Drawing inspiration from the two-phase approach of the TranAD model, our framework elevates this concept by introducing a neural architecture search (NAS) driven methodology. This approach dynamically selects between the refined two-phase approach and our novel iterative self-adversarial approach at each trial. The integration of these advanced techniques addresses the complexities inherent in time series data, particularly in capturing subtle and sophisticated anomalies that traditional encoder-decoder architectures might overlook.
Our development of the iterative self-adversarial approach, in particular, marks a notable enhancement over existing methods. By embedding elements of adversarial training within both the two-phase and iterative approaches, TransNAS-TSAD achieves a comprehensive and adaptable anomaly detection capability. This not only enhances the sensitivity and accuracy of the detection process but also ensures robustness and adaptability across diverse datasets. The synergy of these methodologies within TransNAS-TSAD aligns seamlessly with our research goal of developing a versatile and effective solution for time series anomaly detection, demonstrating our commitment to pushing the boundaries of current technologies.
3.5.1 Two-Phase Adversarial Approach
Drawing from adversarial learning paradigms, the two-phase approach in TransNAS-TSAD seeks to enhance the model’s sensitivity to anomalies through competitive reconstruction stages.
Phase 1 - Preliminary Input Reconstruction
This phase, acting as the foundation, aims for a preliminary reconstruction of the input time-series window. The preliminary input reconstruction yields a focus score, defined by the deviation of the reconstructed output from the actual input:
$L_{\text{focus}}=\|O_{\text{initial}}-W\|^{2}$
where $O_{\text{initial}}$ is the output from the first phase and $W$ is the input time-series window. The resultant “focus score,” derived from the deviations in this initial reconstruction, serves as an attention modulator for the subsequent phase.
Phase 2 - Adversarial, Focus-Driven Reconstruction
Incorporating adversarial elements, this phase utilizes the focus scores from the phase 1 as a guiding beacon. In the adversarial phase, the second decoder aims to maximize the difference between its output and the input. Simultaneously, the first decoder aims to minimize this difference:
$L_{\text{adv1}}=\|O_{\text{adv1}}-W\|^{2}$
$L_{\text{adv2}}=-\|O_{\text{adv2}}-W\|^{2}$
where $O_{\text{adv1}}$ and $O_{\text{adv2}}$ are the outputs from the first and second decoders, respectively, during the adversarial phase.
3.5.2 Iterative Self-Adversarial Approach
Going beyond the structured two-phase reconstruction, TransNAS-TSAD introduces a dynamic iterative approach. This approach, embedded with self-adversarial mechanisms, continually refines its reconstructions.
3.5.3 Iterative Refinement:
Starting with an initial reconstruction, the deviation between the current output and the input provides feedback for the next iteration:
$L_{\text{iteration}}=\left\|O_{\text{current}}-W\right\|_{2}$
The self-adversarial mechanism can be represented as:
$L_{\text{self-adv}}=(L_{\text{iteration, prev}}-L_{\text{iteration, current}})^{2}$
where $L_{\text{iteration, prev}}$ is the loss from the previous iteration and $L_{\text{iteration, current}}$ is the loss from the current iteration. The iterative refinement continues until the change in the loss between consecutive iterations falls below a predetermined threshold, signifying convergence. Mathematically, the convergence criterion can be defined as:
$\Delta L=\left|L_{\text{iteration, current}}-L_{\text{iteration, prev}}\right|$
If $\Delta L<\epsilon$, where $\epsilon$ is a small positive value (e.g., $10^{-5}$), the iteration stops, suggesting that further refinement may not yield significant improvements. After the iterative process converges, the best anomaly score is determined. The best score is derived from the iteration with the smallest reconstruction loss, ensuring that the selected representation most closely matches the input time-series data.
3.5.4 Adaptable and Robust Anomaly Detection
What sets TransNAS-TSAD apart is its adaptability. The model is not confined to a fixed number of reconstruction phases. Depending on the intricacy of the dataset, it can dynamically adjust, ensuring that even the most subtle anomalies are not overlooked. Moreover, the incorporation of adversarial elements, both in the two-phase and iterative approaches, ensures the model continually challenges itself, leading to more refined and accurate reconstructions. TransNAS-TSAD represents a significant advancement in the realm of time-series anomaly detection. By amalgamating auto-regressive inference, adversarial mechanisms, attention modulation, and adaptive iterations, it offers a comprehensive solution, adept at detecting both overt and nuanced anomalies across varied time-series datasets.
3.6 Harnessing the Power of TransNAS-TSAD for Anomaly Detection
Building upon the foundational architecture and adversarial mechanisms of TransNAS-TSAD, we now delve into the practical realm of anomaly detection. This section elucidates the methodologies employed to infer anomalies from time-series data, leveraging the trained transformer model and various augmentation strategies.
3.6.1 The Anomaly Scoring Mechanism
Every incoming data point, denoted as $\hat{W}$, is assessed for anomalies by computing a score that quantitatively reflects its deviation from expected patterns. In this context, $\hat{W}$ represents the new window of the time series data under consideration, and $\hat{C}$ embodies any associated contextual information that could influence the anomaly detection process. The anomaly score, $s$, is calculated as:
$s=\frac{1}{2}\|R_{1}-\hat{W}\|_{2}^{2}+\frac{1}{2}\|\hat{R}_{2}-\hat{W}\|_{2}^{2}$ (7)
where $R_{1}$ and $\hat{R}_{2}$ are the reconstructed outputs from the transformer model, corresponding to different stages of the inference process.
3.6.2 Dual Inference Pathways
TransNAS-TSAD boasts of two distinct inference methodologies:
  • 1.
    Two-Phase Approach: A structured two-step inference process yielding reconstructions $(R_{1},\hat{R_{2}})$.
  • 2.
    Iterative Refinement: For scenarios demanding intricate attention, we engage in iterative inference, perfecting the anomaly scores with each iteration.
Figure 2 illustrates the anomaly detection process on dimension 0 of the SMD test dataset using TransNAS-TSAD. The upper plot displays the true data points (in blue) and the predicted values (in red), while the lower plot indicates the computed anomaly scores and their corresponding labels, offering a visual interpretation of the model’s detection capabilities in identifying anomalies.
Figure 2: Anomaly detection in TransNAS-TSAD on dimension 0 of the SMD test dataset, depicting true versus predicted values and identified anomalies
3.6.3 Dynamic Anomaly Thresholding: The Sentinel
Central to our thresholding strategy is the peaks over threshold (POT) method. At its core, POT establishes a threshold where observations surpassing this mark are deemed anomalous, allowing for effective discernment of extreme data points.
Evolutionary in nature, our modified POT (mPOT) approach ensures adaptability to the ever-changing landscape of time-series data:
$mPOT(x)=POT(x)+\alpha\times\text{recent\_deviation}(x)$ (8)
where $\alpha$ is a weight factor. The function recent_deviation calculates the deviation of the latest data points from their median value. Anomalies are pinpointed whenever any dimension’s score, $s_{i}$, exceeds this dynamic mPOT threshold.
3.6.4 Augmentative Strategies for Enhanced Precision
TransNAS-TSAD employs various augmentative strategies to bolster the precision of the POT method, ensuring that the anomaly detection mechanism remains sensitive, adaptable, and robust. These strategies aim to enhance the detection capabilities by refining the anomaly scores and the thresholds against which they are evaluated.
Moving Average Thresholding (MAT)
MAT is a dynamic thresholding technique that complements the POT method. Instead of using a static threshold, MAT calculates a moving average of recent anomaly scores to adaptively set the threshold. This dynamic adjustment ensures that the threshold is responsive to emerging data trends and patterns, enhancing its relevance and accuracy. The moving average threshold at time $t$ is given by:
$MAT(t)=\frac{1}{N}\sum_{i=t-N}^{t}s_{i}$ (9)
where $(s_{i}$ is the anomaly score at time $i$ and $N$ represents the window size for the moving average.
Rolling Statistics for Nuanced Detection
Rolling statistics, particularly the rolling mean and standard deviation, provide additional context to the POT method by capturing temporal dependencies and trends in the data. These statistics are instrumental in uncovering subtle anomalies that might otherwise be overlooked. For each data point at time $t$, the rolling mean $\mu(t)$ and rolling standard deviation $\sigma(t)$ are computed as:
$\mu(t)=\frac{1}{W}\sum_{i=t-W}^{t}x_{i}\quad\text{and}\quad\sigma(t)=\sqrt{\frac{1}{W}\sum_{i=t-W}^{t}(x_{i}-\mu(t))^{2}}$ (10)
where $x_{i}$ represents the data point at time $i$ and $W$ is the rolling window size. These rolling statistics are incorporated into the feature set, enhancing the model’s ability to discern intricate data variations and improving the relevance of the anomaly scores generated for the POT method.
4 Experimental Setup for TransNAS-TSAD
Our study rigorously evaluates the performance of TransNAS-TSAD against a suite of established benchmark models, including TranAD (Tuli et al., (2022)), LSTM-NDT (Hundman et al., (2018)), DAGMM (Zong et al., (2018)), OmniAnomaly (Su et al., (2019)), MSCRED (Zhang et al., (2019)), MAD-GAN (Li et al., (2019)), USAD (Audibert et al., (2020)), CAE-M (Zhang et al., (2021)), and GDN (Deng & Hooi, (2021)). To ensure a fair comparison, we utilize the hyperparameter configurations as specified in the original publications of these models, relying on their publicly available implementations.
The experimental infrastructure comprises a Google Colab Pro environment, leveraging a NVIDIA Tesla T4 GPU (16GB memory) and an Intel Xeon E5-2670 v3 CPU (8 cores, 51GB memory). The implementation is grounded in Python, utilizing robust libraries such as PyTorch for deep learning, Optuna for hyperparameter optimization, and pandas for data handling. These tools are instrumental in realizing the model implementation, executing the NSGA-II based NAS algorithm, and managing dataset operations.
A pivotal element of our methodology is the neural architecture search (NAS), which automates the architectural design, navigating a comprehensive search space of architectural and training hyperparameters (detailed in Table 1). This search space is meticulously crafted to enable the NAS algorithm to adapt the model architecture to the unique characteristics of each dataset.
Consistent with the practices in OmniAnomaly and TranAD, we apply the peaks over threshold (POT) method (Leadbetter, (1991)) across all datasets with a uniform coefficient of $10^{-4}$. The low quantile parameter, however, is fine-tuned for each specific dataset to maintain alignment with established benchmarks, thus ensuring an equitable basis for comparative analysis.
4.1 Datasets
Dataset Train Size Test Size Dimensions Sequences Anomalies (%)
NAB 4,033 4,033 1 6 0.92
UCR 1,600 5,900 1 4 1.88
MBA 100,000 100,000 2 8 0.14
SMAP 135,183 427,617 25 55 13.13
MSL 58,317 73,729 55 3 10.72
SWaT 496,800 449,919 51 1 11.98
WADI 1,048,571 172,801 123 1 5.99
SMD 708,405 708,420 38 4 4.16
Table 2: Dataset Statistics
Our study employs a comprehensive suite of datasets, each meticulously chosen to challenge and validate the robustness of time series anomaly detection methods across a variety of real-world scenarios. Table 2 summarizes the statistics of the datasets. These datasets, renowned for their complexity and diversity, allow for an effective benchmarking of our TransNAS-TSAD approach against existing methodologies.
  • 1.
    Numenta Anomaly Benchmark (NAB) Dataset: OThis dataset encompasses a wide range of real-world data, including temperature sensor readings, cloud machine CPU utilization, service request latencies, and taxi demand data in New York City. It’s important to note some anomalies in labeling within this dataset, particularly in the nyc-taxi trace, which is often excluded from analyses (Ahmad et al., (2017); Nakamura et al., (2020)).
  • 2.
    HexagonML (UCR) Dataset: Featured in the KDD 2021 cup, this dataset comprises a diverse collection of univariate time series. In our research, we selectively utilized datasets that are derived from real-world sources, focusing specifically on the InternalBleeding and ECG datasets. This choice was driven by our need for natural data representations in our analysis. Consequently, we did not include any synthetic sequences that are also part of the HexagonML (UCR) Dataset (Dau et al., (2018)).”
  • 3.
    MIT-BIH Supraventricular Arrhythmia Database (MBA) Dataset: This dataset includes electrocardiogram recordings from four patients, featuring anomalies such as supraventricular contractions or premature heartbeats. It’s a well-recognized dataset in data management studies (Moody & Mark, (2001); Boniol et al., (2020)).
  • 4.
    Soil Moisture Active Passive (SMAP) Dataset:: Provided by NASA, this dataset consists of soil samples and telemetry data from the Mars rover (Hundman et al., (2018)).
  • 5.
    Mars Science Laboratory (MSL) Dataset: Similar to the SMAP dataset, the MSL dataset includes sensor and actuator data from the Mars rover.(Hundman et al., (2018)) Due to the presence of many trivial sequences, only specific non-trivial sequences are typically analyzed such as (A4, C2 and T1) pointed by (Nakamura et al., (2020); Tuli et al., (2022)).
  • 6.
    Secure Water Treatment (SWaT) Dataset: This dataset is derived from a real-world water treatment plant, including data from 7 days of normal operations and 4 days under abnormal conditions. It features readings from various sensors and actuators (Mathur & Tippenhauer, (2016)).
  • 7.
    Water Distribution (WADI) Dataset: An expansion of the SWaT dataset, WADI includes a larger array of sensors and actuators. The dataset spans over a longer period, with 14 days of normal operation and 2 days under attack scenarios (Ahmed et al., (2017)).
  • 8.
    Server Machine Dataset (SMD): Covering five weeks of data, this dataset includes resource utilization traces from 28 machines in a compute cluster. Only specific non-trivial sequences are used for analysis (Su et al., (2019)).
Each dataset is selected for its potential to challenge and validate the robustness of time series anomaly detection methods across various industry scenarios.
4.2 Evaluation Criteria
In our study, aimed at advancing anomaly detection in time series data, we employ key metrics aligned with the objectives of the TransNAS-TSAD framework. These metrics are selected to optimize anomaly detection effectiveness, considering practical deployment aspects.
4.2.1 F1 Score as Primary Performance Indicator
The F1 score, the harmonic mean of precision and recall, is central to our evaluation strategy. It is particularly effective in addressing the challenge of imbalanced datasets common in anomaly detection, offering a balanced performance measure by accounting for the trade-off between false positives and false negatives.
4.2.2 Introducing the Efficiency-Accuracy-Complexity Score (EACS)
Model complexity is often indicated by parameter count. However, this metric can be misleading, especially for transformer-based models like TransNAS-TSAD. These models, despite their high parameter count, capture complex data patterns effectively, justifying their complexity. Therefore, we report parameter count and interpret it in the context of the model’s advanced capabilities. To comprehensively evaluate TransNAS-TSAD, we introduce the Efficiency-Accuracy-Complexity Score (EACS), a composite metric encompassing accuracy, training efficiency, and model complexity. It is defined as:
$EACS=w_{\text{acc}}\times\left(\frac{\text{F1 Score}}{\text{Max F1 Score}}\right)+w_{\text{time}}\times\left(\frac{\text{Training Time}}{\text{Max Training Time}}\right)+w_{\text{params}}\times\left(\frac{\text{Parameter Count}}{\text{Max Parameter Count}}\right)$
where $w_{acc}$, $w_{time}$, and $w_{params}$ are weights for the normalized F1 score, inverse training time, and inverse parameter count, respectively. Here, the weights are set as $w_{acc}=0.4$, $w_{time}=0.4$, and $w_{params}=0.2$, balancing accuracy, training efficiency, and model simplicity. This weighting scheme aligns with our research goals and practical deployment considerations, prioritizing accuracy and training efficiency while considering model complexity.
Transformer models, known for their high parameter count, offer efficient processing capabilities through parallelization, making them suitable for scenarios requiring high accuracy where computational resources are not a constraint.
4.3 Note on Additional Metrics
Other metrics, such as ROC/AUC, are also valuable; however, our focus on the F1 score, parameter count, and EACS aligns with the specific goals of TransNAS-TSAD. This approach ensures a balanced evaluation of the models, considering both their effectiveness in anomaly detection and practical deployability.
5 Results
5.1 Evaluating TransNAS-TSAD: Precision, Recall, and F1 Score Analysis
In our detailed evaluation of anomaly detection models, as captured in Table 3, the TransNAS-TSAD model emerges as a formidable contender, showcasing its prowess across diverse datasets including NAB, UCR, MBA, SMAP, MSL, SWaT, WADI, and SMD, with metrics of precision, recall, and F1 scores serving as the benchmarks.
Standing out for its exceptional performance, TransNAS-TSAD notably surpasses other models in datasets such as NAB, UCR, MBA, and SMAP, with remarkable F1 scores (0.9411, 0.9910, 0.9861, and 0.9510 respectively), thereby demonstrating its adeptness in handling both univariate and multivariate time series. This is particularly evident when compared to models like OmniAnomaly and MSCRED, where TransNAS-TSAD’s ability to navigate complex data patterns becomes apparent.
However, in the MSL dataset, the TranAD model, with its nuanced approach, shows a marginally superior performance with an F1 score of 0.9494 compared to TransNAS-TSAD’s 0.9482, possibly indicative of TranAD’s specific strengths in handling the dataset’s characteristics. Despite this, TransNAS-TSAD regains its lead in the SWaT dataset with an F1 score of 0.8314, highlighting its adaptability to various industrial contexts.
Its exemplary performance continues in the challenging WADI and SMD datasets, where it achieves the highest F1 scores of 0.7735 and 0.9986, respectively, showcasing its unrivaled efficiency in detecting subtle anomalies in complex, high-dimensional datasets. This overarching success of TransNAS-TSAD, especially when contrasted with TranAD’s performance, underscores its comprehensive capability in anomaly detection.
While models like LSTM-NDT and DAGMM show variable performance across different datasets, their inconsistency underlines TransNAS-TSAD’s steady effectiveness across various scenarios. Similarly, while OmniAnomaly and MSCRED perform well in certain contexts, they fall short of TransNAS-TSAD’s sophisticated temporal pattern handling.
Other competitive models, including MAD-GAN and USAD, are eclipsed by TransNAS-TSAD’s more holistic approach, and even though CAE-M and GDN demonstrate strengths in specific scenarios, they generally do not match TransNAS-TSAD’s overall adaptability.
The superiority of TransNAS-TSAD, especially in comparison to TranAD, invites further investigation through critical difference analysis, which would reinforce the statistical significance of the observed performance differences and provide a deeper understanding of TransNAS-TSAD’s robustness and reliability against a backdrop of diverse and challenging anomaly detection scenarios.
Method NAB UCR MBA SMAP
P R F1 P R F1 P R F1 P R F1
LSTM-NDT 0.6400 0.6667 0.6531 0.5231 0.8294 0.5231 0.9207 0.9718 0.9456 0.8523 0.7326 0.7879
DAGMM 0.7622 0.7292 0.7453 0.5337 0.9718 0.5337 0.9475 0.9900 0.9683 0.8069 0.9891 0.8888
OmniAnomaly 0.8421 0.6667 0.7442 0.8346 0.9999 0.8346 0.8561 1.000 0.9225 0.813 0.9419 0.8728
MSCRED 0.8522 0.6700 0.7502 0.5441 0.9718 0.5441 0.9272 1.0000 0.9623 0.8175 0.9216 0.8664
MAD-GAN 0.8666 0.7012 0.7752 0.8538 0.9891 0.8538 0.9396 1.0000 0.9689 0.8157 0.9216 0.8654
USAD 0.8421 0.6667 0.7442 0.8952 1.0000 0.8952 0.8953 0.9989 0.9443 0.748 0.9627 0.8419
CAE-M 0.7918 0.8019 0.7968 0.6981 1.0000 0.6981 0.8442 0.9997 0.9154 0.8193 0.9567 0.8827
GDN 0.8129 0.7872 0.7998 0.6894 0.9988 0.6894 0.8832 0.9892 0.9332 0.748 0.9891 0.8518
TranAD 0.8889 0.9892 0.9364 0.9407 1.0000 0.9407 0.9569 1.0000 0.978 0.8043 0.9999 0.8915
TransNAS-TSAD 0.8888 1.0000 0.9411 0.9823 1.0000 0.9910 0.9726 1.0000 0.9861 0.9066 1.0000 0.9510
Method MSL SWaT WADI SMD
P R F1 P R F1 P R F1 P R F1
LSTM-NDT 0.6288 1.0000 0.7721 0.7778 0.5109 0.6167 0.0138 0.7823 0.0271 0.9736 0.844 0.9042
DAGMM 0.7363 1.0000 0.8482 0.9933 0.6879 0.8128 0.076 0.9981 0.1412 0.9103 0.9914 0.9491
OmniAnomaly 0.7848 0.9924 0.8765 0.9782 0.6957 0.8131 0.3158 0.6541 0.426 0.8881 0.9985 0.9401
MSCRED 0.8912 0.9862 0.9363 0.9992 0.677 0.8072 0.2513 0.7319 0.3741 0.7276 0.9974 0.8414
MAD-GAN 0.8516 0.993 0.9169 0.9593 0.6957 0.8065 0.2233 0.9124 0.3588 0.9991 0.844 0.915
USAD 0.7949 0.9912 0.8822 0.9977 0.6879 0.8143 0.1873 0.8296 0.3056 0.906 0.9974 0.9495
CAE-M 0.7751 1.0000 0.8733 0.9697 0.6957 0.8101 0.2782 0.7918 0.4117 0.9082 0.9671 0.9367
GDN 0.9308 0.9892 0.9591 0.9697 0.6957 0.8101 0.2912 0.7931 0.426 0.717 0.9974 0.8342
TranAD 0.9038 0.9999 0.9494 0.976 0.6997 0.8151 0.3529 0.8296 0.4951 0.9262 0.9974 0.9605
TransNAS-TSAD 0.9567 1.000 0.94.82 0.9415 0.7624 0.8314 0.8626 0.7340 0.7735 0.9985 0.9988 0.9986
Table 3: Benchmarking TransNAS-TSAD against Baseline Models: A summary of precision (P), recall (R), and F1 scores showcases the advanced anomaly detection capabilities of TransNAS-TSAD, derived via NSGA-II-based NAS, across diverse datasets.
5.2 In-Depth Analysis of Training Efficiency and EACS in Anomaly Detection Models
Furthering our investigation from Table 3, Table 4 provides a more granular look at the training efficiency and overall performance of anomaly detection models, with a particular emphasis on the Efficiency-Accuracy-Complexity Score (EACS). This table not only reiterates the superior performance of TransNAS-TSAD in precision and recall metrics, as observed in Table 3 but also introduces a critical analysis of training times and parameter counts across a range of datasets including NAB, UCR, MBA, SMAP, MSL, SWaT, WADI, and SMD.
The training time, a pivotal yet often overlooked aspect, emerges as a decisive factor in evaluating the practicality of models. In operational settings, where rapid model deployment and iterative updating are essential, training time gains precedence over traditional metrics like parameter count. This is evidenced in the case of TransNAS-TSAD, which, across diverse datasets, demonstrates not only high F1 scores but also markedly reduced training times, leading to its high EACS. In the NAB dataset, for instance, TransNAS-TSAD achieves an F1 score of 94.11% with a training time of just 2.70 sec, resulting in an EACS of 0.9742. This trend of balancing accuracy with training efficiency is replicated in the UCR dataset, where it achieves an F1 score of 99.10% and an even lower training time of 2.24 sec, leading to an EACS of 0.9922.
Diving deeper, in datasets like MBA and SMAP, TransNAS-TSAD’s EACS scores of 0.9932 and 0.9734, respectively, are not solely the product of its accuracy but are significantly bolstered by its training time efficiency. For example, in MBA, its training time of 2.27 is among the lowest, demonstrating the model’s capability to quickly adapt and learn from complex multivariate time series data. In comparison, models such as MSCRED, while exhibiting commendable F1 scores, are penalized in their EACS due to their disproportionately high training times (e.g., 774.99 sec for MBA), highlighting a crucial trade-off between model complexity and operational efficiency.
In more challenging datasets like SWaT and WADI, the importance of training time becomes even more pronounced. TransNAS-TSAD’s ability to maintain high F1 scores while drastically cutting down on training times positions it as a highly effective tool for real-world applications. Its EACS of 0.9297 for SWaT and 0.6977 for WADI, while highlighting its proficiency in these industrial datasets, also reflect the balanced approach of the model, juxtaposing accuracy with swift adaptability.
Furthermore, the analysis of models like GDN and TranAD provides additional context. While GDN shows strong performance in specific datasets like MSL with the highest F1 score (95.91%) and a respectable EACS of 0.9191, it falls behind TransNAS-TSAD in terms of training efficiency. Similarly, TranAD, despite its high F1 scores and relatively lower parameter counts, does not consistently match the training efficiency of TransNAS-TSAD, as seen in datasets like SMD and WADI.
In conclusion, Table 4’s comprehensive analysis underscores the critical role of training time in the practical deployment of anomaly detection models. It establishes that models with lower training times, like TransNAS-TSAD, offer a significant advantage in real-world scenarios, effectively balancing accuracy, efficiency, and complexity. This detailed examination not only affirms the superiority of TransNAS-TSAD in standard performance metrics but also in its operational readiness, making it an exemplary model for efficient and effective anomaly detection in diverse and demanding environments.
Method NAB UCR MBA SMAP
F1 Training Time # Params EACS F1 Training Time # Params EACS F1 Training Time # Params EACS F1 Training Time # Params EACS
LSTM-NDT 65.31 23.40 1710 0.8424 52.31 11.14 1765 0.7985 94.56 47.80 18641 0.9520 78.79 36.43 29146 0.8459
DAGMM 74.53 64.50 1266 0.8467 53.37 27.50 1214 0.7919 96.83 92.36 2448 0.9394 88.80 26.60 7266 0.9050
OmniAnomaly 74.42 88.70 13717 0.8250 83.46 38.21 13025 0.8863 92.25 197.66 25474 0.8649 87.28 66.77 16813 0.8230
MSCRED 75.02 503.60 1237377 0.3001 54.41 559.60 128655 0.2176 96.23 774.99 2441778 0.3849 86.64 25.70 8237452 0.6982
MAD-GAN 77.52 53.90 838 0.8671 85.38 31.50 822 0.9177 96.89 229.78 1877 0.8688 86.54 41.26 6718 0.8683
USAD 74.42 65.40 1359 0.8455 89.52 54.23 1359 0.9172 94.43 191.26 1609 0.8789 84.19 49.22 7395 0.8439
CAE-M 79.68 33.60 7229 0.8909 69.81 22.10 7365 0.8520 91.54 89.66 15411 0.9186 88.27 212.51 7229 0.5529
GDN 79.98 131.60 574 0.8153 68.94 64.70 566 0.8286 93.32 203.45 1106 0.8682 85.18 104.50 2974 0.7440
TranAD 93.64 3.37 615 0.9718 94.07 1.14 619 0.9745 97.80 5.11 1298 0.9885 89.15 5.62 62271 0.9445
TransNAS-TSAD 94.11 2.70 285 0.9742 99.10 2.24 1690 0.9922 98.61 2.27 593 0.9932 94.60 3.19 39045 0.9734
Method MSL SWaT WADI SMD
F1 Training Time # Params EACS F1 Training Time # Params EACS F1 Training Time # Params EACS F1 Training Time # Params EACS
LSTM-NDT 77.21 37.80 61856 0.8887 66.70 41.50 72854 0.7957 2.71 422.67 133712 0.5737 90.42 460.70 42853 0.9100
DAGMM 84.82 25.66 18756 0.9258 81.28 29.66 16558 0.8748 14.12 246.32 39387 0.6392 94.91 337.90 10516 0.9420
OmniAnomaly 87.65 29.69 39753 0.9348 81.31 40.20 36541 0.8568 42.60 361.60 39753 0.7471 94.01 311.20 22941 0.9412
MSCRED 93.63 40.70 18476728 0.7535 80.72 236.80 14227264 0.3229 37.41 1884.25 75530 0.6435 84.14 349.77 14237356 0.6978
MAD-GAN 91.69 32.64 17446 0.9497 80.65 34.70 15958 0.8638 35.88 321.20 35764 0.7228 91.50 424.60 9903 0.9188
USAD 88.22 29.60 14859 0.9374 81.43 35.81 12502 0.8651 30.56 389.72 32859 0.6983 94.95 328.77 10609 0.9432
CAE-M 87.33 774.60 204687 0.5471 81.01 71.22 194525 0.8010 41.17 7724.92 388905 0.3203 93.67 3606.90 114219 0.5731
GDN 95.91 121.80 153541 0.9191 81.01 83.55 115442 0.7813 42.60 6047.12 287082 0.4245 83.42 1000.40 7255 0.8226
TranAD 94.94 6.88 272181 0.9733 81.53 2.10 31336 0.9221 49.51 177.60 1378173 0.6316 96.05 56.70 135110 0.9760
TransNAS-TSAD 94.82 5.44 236491 0.9739 83.14 1.60 13464 0.9297 77.35 225.60 1752919 0.6977 99.81 52.40 132050 0.9916
Table 4: Comparative analysis of the Efficiency-Accuracy-Complexity Score (EACS) across various anomaly detection models, highlighting the performance of TransNAS-TSAD. The EACS is calculated by normalizing and weighting F1 scores (0.4), training times (0.4), and parameter counts (0.2), demonstrating the balanced efficiency and accuracy of TransNAS-TSAD against benchmark models across multiple datasets
Arch and Training Params SMAP UCR MBA SWaT MSL SMD NAB WADI
Learning Rate 0.0002128 0.0019997 0.003441 0.0000403 0.006192 0.0002273 0.006924 0.0000131
Dropout Rate 0.2353 0.4474 0.1795 0.3836 0.3730 0.1554 0.4560 0.2237
Dim Feedforward 101 124 41 42 121 86 10 16
Batch Size 32 48 16 32 32 48 96 112
Encoder Layers 2 1 3 2 2 3 1 2
Decoder Layers 1 2 1 3 3 1 2 1
Activation Func. sigmoid leaky_relu tanh tanh leaky_relu sigmoid relu sigmoid
Time Warping False True True True True False False False
Time Masking True True False True True False True False
Gaussian Noise 0.000151 0.027956 0.007925 0.058019 0.000428 0.000628 0.000119 0.000194
Linear Embedding False True False False True True True True
Phase Type 2phase iterative iterative 2phase 2phase 2phase iterative 2Phase
Self Conditioning False True False True True False True True
Layer Norm False False False True True True False True
Pos. Enc. Type sinusoidal fourier fourier sinusoidal fourier sinusoidal sinusoidal fourier
FFN Layers 1 1 1 3 2 3 1 3
Attn Heads 25 1 2 51 55 38 1 127
Window Size 10 20 14 22 26 12 18 26
Table 5: TransNAS-TSAD Hyperparameter Optimization Results: Detailed summary of the optimal training and architectural hyperparameters identified for each dataset, demonstrating the adaptive precision of TransNAS-TSAD in time series anomaly detection.
5.2.1 Optimal Model Configurations Identified by TransNAS-TSAD
Table 5 showcases the culmination of TransNAS-TSAD’s NSGA-II optimization process, highlighting the training and architectural hyperparameters of models that achieved the best balance between high F1 scores and low parameter counts. These configurations represent the pinnacle of the framework’s capability to identify models uniquely suited to each dataset’s anomaly detection needs.
A key observation from the table is the diversity in architectural parameters across datasets, reflecting TransNAS-TSAD’s adeptness in tailoring models to specific data characteristics. For example, in the SMAP dataset, the model employs a relatively simple architecture with fewer encoder and decoder layers, which is effective in handling the environmental time series data of this dataset. In contrast, the WADI dataset, known for its complex sensor network, necessitates a more intricate model structure, evident in its higher number of attention heads and the use of a two-phase reconstruction and refinement strategy.
Similarly, the variation in window sizes, ranging from 12 data points for SMD to 30 data points for WADI, underscores the model’s flexibility in adapting to the temporal scale of different datasets. Larger window sizes in datasets like WADI allow the model to capture longer-term dependencies and subtle anomalies in extensive time series data, a crucial requirement for sophisticated water-related infrastructures.
The adaptability of TransNAS-TSAD is further demonstrated in its choice of positional encoding and layer normalization techniques, which vary significantly among datasets. For instance, the sinusoidal positional encoding in SMAP and NAB caters to their unique temporal patterns, while the Fourier positional encoding in UCR and WADI aligns with the datasets’ complex spectral characteristics.
These results from TransNAS-TSAD’s NAS process validate the framework’s ability to not only fine-tune hyperparameters for operational efficiency but also to intricately adapt its architectural design to the nuanced demands of various time series anomaly detection scenarios, ensuring optimal detection performance across a broad spectrum of datasets.
Figure 3: Analysis of hyperparameter importance for F1 score optimization in TransNAS-TSAD across four datasets (NAB, MBA, SMAP, and WADI). The plots illustrate the relative impact of different hyperparameters on the F1 score, guiding the model’s fine-tuning process for effective anomaly detection in diverse time series datasets
5.3 Architectural and Training Hyperparameter Importance for F1 Score Optimization
The optimization of F1 scores in the TransNAS-TSAD model shown in Figure 3 is critically analyzed through hyperparameter importance plots for four datasets: NAB, MBA, SMAP, and WADI. These plots offer quantitative insights into the relative impact of various hyperparameters on the F1 score, aiding the fine-tuning process in our NAS-driven anomaly detection model.
For the NAB dataset, gaussian_noise and activation_function are the most influential hyperparameters, with importances of 0.18 and 0.15, respectively. The incorporation of gaussian_noise enhances the model’s generalization capabilities, essential for robustness, while the choice of activation_function is key for capturing the data’s nonlinear relationships.
The MBA dataset’s analysis underscores the importance of activation_function as 0.11, num_feedforward_layers as 0.07, and dropout_rate as 0.07, indicating the need for a complex model architecture to effectively learn from multivariate ECG time series.
In the SMAP dataset, gaussian_noise with importance 0.53 is paramount, highlighting the model’s need to handle noisy environmental data effectively. Additionally, batch_size with importance 0.07 influences the model’s performance, reflecting the impact of batch processing on training.
The WADI dataset, with its extensive sensor network, prioritizes activation_function with importance 0.16 and dropout_rate with importance 0.08 to maintain model robustness and prevent overfitting in high-dimensional spaces.
The recurrence of activation_function and dropout_rate across datasets emphasizes their role in non-linear data transformation and regularization. This pattern reflects the impact of data quality and structure on the model’s learning efficacy.
The conducted hyperparameter analysis illuminates the tailored performance of TransNAS-TSAD across varied datasets, with each dataset’s specifics dictating the importance of particular hyperparameters. This not only affirms the model’s adaptability but also its capacity for ongoing enhancement, ensuring it remains adept at confronting the evolving complexities of time series data.
Figure 4: Pareto front plots illustrating the trade-off between F1 score and number of parameters for NAB, MBA, SMAP, and WADI datasets in the TransNAS-TSAD optimization process.
5.4 Pareto Front Analysis for Model Optimization
Building upon the insights from the hyperparameter importance analysis, the Pareto front optimization results for the TransNAS-TSAD model, as depicted in Figure 4, further elucidate the model’s performance. The Pareto front plots for the NAB, MBA, SMAP, and WADI datasets underscore the delicate balance between model complexity, as measured by the number of parameters, and the model’s effectiveness, as quantified by the F1 score.
For the NAB dataset, the Pareto front indicates a dense congregation of model configurations. A significant number of these configurations exhibit a high parameter count without a corresponding increase in the F1 score, suggesting a potential plateau in performance gains relative to complexity. This observation prompts a critical evaluation of model parsimony, encouraging the selection of simpler models that maintain performance while mitigating the risk of overfitting.
In the MBA dataset, the dispersion of trials across the Pareto front reflects a comprehensive exploration of the architectural space. Interestingly, several models achieve commendable F1 scores without a proportionate surge in parameters, highlighting efficient architectural choices that capture the essential dynamics of ECG time series data without unnecessary complexity.
The SMAP dataset presents an outlier model with a substantial number of parameters yet achieving a high F1 score. This result may point to an overfitting scenario where the model’s complexity does not translate into generalized performance. Such an insight is invaluable for guiding the model selection process towards architectures that balance accuracy with generalization.
The Pareto front for the WADI dataset, characterized by its complex sensor network, illustrates the necessity of sophisticated models to enhance the F1 score. The trend of increasing model complexity to achieve incremental improvements in performance is evident, underlining the inherent challenges of anomaly detection in high-dimensional industrial control systems.
These Pareto fronts provide a visual and quantitative tool for identifying optimal model configurations that achieve a balance between performance and complexity. They serve as a decision-making aid for selecting models that not only excel in accuracy but are also efficient and scalable, aligning with the practical demands of anomaly detection in diverse environments.
The findings from the Pareto front analysis are integral to the ongoing development of the TransNAS-TSAD model within the broader scope of our research. They contribute to the understanding of how different model architectures perform across various datasets, offering a pathway to enhancing the model’s adaptability and ensuring its continued efficacy in the face of evolving data challenges.
6 Discussion: Addressing Challenges and Exploring Future Directions in TransNAS-TSAD
TransNAS-TSAD represents a significant advancement in anomaly detection within time series data, primarily through its innovative integration of advanced adversarial learning paradigms, NSGA-II optimization, and transformer architecture optimization. This confluence of technologies marks a substantial leap forward, particularly in the detection of subtle and complex anomalies that have typically eluded traditional methods.
A pivotal challenge addressed by TransNAS-TSAD is the balance of detection accuracy against computational efficiency. The NSGA-II algorithm plays a crucial role in this regard, effectively managing the equilibrium to ensure optimal model performance, without incurring excessive computational demands. This aspect is particularly crucial in real-world applications where resources are finite and efficiency is paramount.
The adaptability of TransNAS-TSAD is further exemplified in its dynamic adjustment capabilities, allowing it to effectively respond to the unique characteristics of different datasets. This adaptability is essential in the domain of anomaly detection, where dataset variability can present diverse challenges. Additionally, the iterative self-adversarial approach employed by TransNAS-TSAD significantly enhances detection accuracy, showcasing the model’s sophisticated capabilities in identifying anomalies.
However, challenges remain, particularly in the realm of ensuring model generalization across a diverse array of datasets. The Pareto front analysis within TransNAS-TSAD has highlighted a delicate balance between model complexity and effectiveness. Some configurations risk overfitting, which is a pertinent issue for future research endeavors. Improving generalization capabilities, without compromising detection accuracy, remains a key area for further investigation.
Looking ahead, several promising avenues for enhancement and innovation present themselves. Enhanced real-time data processing capabilities, particularly for applications in environmental monitoring and industrial control systems, represent a significant area for advancement. Techniques from data assimilation and online learning could be effectively integrated into TransNAS-TSAD to address these challenges. Additionally, the development and implementation of dynamic thresholding strategies, such as the Moving Average Thresholding (MAT) and the incorporation of rolling statistics, offer exciting prospects. These augmentative strategies could pave the way for even more nuanced detection capabilities within the TransNAS-TSAD framework.
Furthermore, the exploration of hybrid systems that synergize machine learning with simulation approaches, along with advancements in neuro-symbolic systems, could substantially enhance the model’s adaptability and effectiveness across various scenarios. Finally, a human-centric approach to machine learning, integrating human feedback in a more intuitive and formalized manner, remains a significant challenge. TransNAS-TSAD stands to benefit greatly from such integration, enhancing not only the interpretability but also the overall usability of the model in real-world applications.
As the landscape of time series anomaly detection continues to evolve, so too will the strategies and methodologies embodied in TransNAS-TSAD, ensuring its continued relevance and efficacy in this dynamic field.
7 Conclusion
TransNAS-TSAD emerges as a groundbreaking framework in the realm of time series anomaly detection, effectively integrating transformer architecture with neural architecture search and NSGA-II optimization. The model’s superior performance, evidenced across diverse datasets, demonstrates its robustness in accurately detecting anomalies in both univariate and multivariate time series data.
The strategic application of NSGA-II within TransNAS-TSAD skillfully addresses the challenge of balancing accuracy with computational efficiency, a critical consideration in practical applications. Moreover, the incorporation of advanced adversarial learning paradigms contributes to its ability to detect nuanced anomalies, a notable advancement over traditional methods.
TransNAS-TSAD’s development signifies a notable contribution to the field of anomaly detection, enhancing the capabilities and efficiency of current methodologies. As the landscape of data analysis continues to evolve, the principles and approaches embodied in TransNAS-TSAD will undoubtedly inspire further advancements in machine learning, paving the way for innovative applications in various sectors. The framework not only sets a new benchmark in anomaly detection but also opens up exciting prospects for future research, particularly in enhancing real-time processing and integrating human-centric approaches to machine learning.
Acknowledgments
This material is based upon work supported by the National Science Foundation under Grant No. EAR 2012123. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Any use of trade, firm, or product names is for descriptive purposes only and does not imply endorsement by the U.S. Government. The work was also supported by the University of Vermont College of Engineering and Mathematical Sciences through the REU program. We would like to extend our gratitude to the providers of the data sets used in this work, whose contributions were invaluable to our research. The authors would like to thank the US Geological Survey (USGS) for offering the domain expertise that was crucial to identify the peak anomaly types that are of practical importance.

### 04 Mar 2024 v3

@@
-TransNAS-TSAD: Harnessing Transformers for Multi-Objective Neural Architecture Search in Time Series Anomaly Detection [^1] This document is the results of the research project funded by the National Science Foundation.22footnotemark: 2
+TransNAS-TSAD: Harnessing Transformers for Multi-Objective Neural Architecture Search in Time Series Anomaly Detection
 Abstract
 The surge in real-time data collection across various industries has underscored the need for advanced anomaly detection in both univariate and multivariate time series data.
-Traditional methods, while comprehensive, often struggle to capture the complex interdependencies in such data.
-This paper introduces TransNAS-TSAD, a novel framework that synergizes transformer architecture with neural architecture search (NAS), enhanced through NSGA-II algorithm optimization.
-This innovative approach effectively tackles the complexities of both univariate and multivariate time series, balancing computational efficiency with detection accuracy.
-Our evaluation reveals that TransNAS-TSAD surpasses conventional anomaly detection models, demonstrating marked improvements in diverse data scenarios.
-We also propose the Efficiency-Accuracy-Complexity Score (EACS) as a new metric for assessing model performance, emphasizing the crucial balance between accuracy and computational resources.
+This paper introduces TransNAS-TSAD, a framework that synergizes the transformer architecture with neural architecture search (NAS), enhanced through NSGA-II algorithm optimization.
+This approach effectively tackles the complexities of time series data, balancing computational efficiency with detection accuracy.
+Our evaluation reveals that TransNAS-TSAD surpasses conventional anomaly detection models due to its tailored architectural adaptability and the efficient exploration of complex search spaces, leading to marked improvements in diverse data scenarios.
+We also introduce the Efficiency-Accuracy-Complexity Score (EACS) as a new metric for assessing model performance, emphasizing the balance between accuracy and computational resources.
 TransNAS-TSAD sets a new benchmark in time series anomaly detection, offering a versatile, efficient solution for complex real-world applications.
-This research paves the way for future developments in the field, highlighting its potential in a wide range of industry applications.
+This research highlights the TransNAS-TSAD potential across a wide range of industry applications and paves the way for future developments in the field.
 keywords:
-Time Series Anomaly Detection, Transformers in Machine Learning, Neural Architecture Search (NAS), NSGA-II Optimization, Deep Learning for Time Series Analysis, Adversarial Learning Paradigms
-[…]
-[1]organization=Department of Computer Science, addressline=University of Vermont,postcode=05405,city=Burlington, VT, country=U.S.A
+Neural Architecture Search, Time Series Anomaly Detection, Transformers, Multi-Objective Optimization
+ [^1] The source code, data, and/or other artifacts have been made available at https://github.com/ejokhan/TransNAS_TSAD.git.
 1 Introduction
-The prevalence of time series data in various sectors—from finance to healthcare, from infrastructure to manufacturing—highlights its critical role in modern analytics.
-Such data are often pivotal in uncovering underlying patterns and dependencies that are vital for strategic insights and anomaly detection, which can indicate significant shifts in system behaviors or emerging issues (Brophy et al., (2023); Li et al., (2023)).
-Traditional statistical methods have long been the mainstay for anomaly detection in time series data, trusted for their robust mathematical frameworks (Thudumu et al., (2020); Wang et al., (2011); Leadbetter, (1991)).
-Yet, the advent of Big Data—with its hallmark features of volume, velocity, and variety—has stretched these methods to their limits, revealing shortcomings through high false positive rates and missed detections (Ashabi et al., (2020); Al-Sai et al., (2019)).
-In contrast, deep learning has ushered in a paradigm shift, offering enhanced adaptability and performance in anomaly detection (Ismail Fawaz et al., (2019); Ma et al., (2021); Haq et al., (2023)).
-The transformer architecture, in particular, has emerged as a groundbreaking development due to its self-attention mechanism, enabling significant strides not only in natural language processing but also in time series analysis (Vaswani et al., (2017); Tuli et al., (2022); Kim et al., (2023)).
-However, the specific application of transformer models to time series anomaly detection necessitates further exploration to fine-tune their structures and parameters for optimal performance (Arslan et al., (2023)).
-Our initiative, TransNAS-TSAD, represents a significant shift from traditional transformer models, such as TranAD, to a more dynamic and adaptable framework.
-This new model integrates the transformer’s architectural strengths with systematic optimization through neural architecture search (NAS), emphasizing the multi-objective optimization strengths of the NSGA-II algorithm (Elsken et al., (2019); Deb et al., (2002)).
-TransNAS-TSAD marks a significant evolution from TranAD’s static architecture, primarily through the integration of NSGA-II.
-This shift, grounded in (Lu et al., (2023)) study, leverages NSGA-II’s proven superiority in exploring complex search spaces more efficiently than recent algorithms.
-By adeptly balancing performance with computational efficiency, NSGA-II enhances our neural architecture design for anomaly detection, offering improved adaptability and efficiency.
-(Lu et al., (2020).
-1.1 Research Context and Challenges
-Adapting anomaly detection models for time series data is a complex task, especially when dealing with multivariate series that present high-dimensional challenges and necessitate sophisticated model architectures (Wei, (2018); Liu et al., (2018)).
-The selective attention capabilities of the transformer model present a compelling solution, but only if the architectures are tailored to the specific characteristics of the dataset (Lian et al., (2019)).
-The marriage of NAS with multi-objective optimization algorithms like NSGA-II represents a strategic evolution in automated model design, offering a methodical path to discovering optimal architectures that consider multiple, often conflicting, objectives (Liu et al., (2021); Xue et al., (2023)).
-1.2 Theoretical Foundations and Methodological Approach
-Our methodology builds on the interplay between the theoretical principles of evolutionary algorithms and the empirical strengths of deep learning architectures.
-We aim to craft a robust NAS framework that can intelligently navigate and select transformer architectures with hyperparameters fine-tuned for anomaly detection in time series data (Wang et al., (2021); Ma et al., (2021)).
-The use of NSGA-II is central to this approach, providing a mechanism to balance the trade-offs between detection accuracy and model complexity (Wang et al., (2021)).
-1.3 Guiding Philosophies and Contributions
-Our research is guided by the following key principles:
-  • 1.
-    Adaptation to Time Series Data: We assert the need for specialized models that respect the unique temporal and dimensional characteristics of time series data.
-  • 2.
-    Maximizing Transformer Utility: We aim to exploit the full capabilities of the transformer architecture, particularly for temporal data interpretation.
-  • 3.
-    Strategic Optimization via NAS: We employ NAS, especially NSGA-II, as a systematic tool for refining architectures towards optimal model configurations.
-Our contributions are as follows:
-  1. 1.
-    Novel Application of NAS: We leverage NAS to uniquely optimize transformer architectures for the domain of time series anomaly detection.
-  2. 2.
-    Architectural Parameter Exploration: We conduct extensive experimentation to identify transformer configurations that redefine design norms and improve detection performance.
-  3. 3.
-    Benchmarking Performance: We set new standards for performance, outperforming existing methods in a variety of datasets.
-  4. 4.
-    Model Design Efficiency: Our approach yields models that balance computational efficiency with analytical accuracy.
-  5. 5.
-    Architectural and Search Strategy Synergy: We seamlessly blend the structural advantages of transformers with the adaptive capabilities of NAS, resulting in models adept for time series anomaly detection.
+The ubiquity of time series data across various sectors, ranging from finance [10] and healthcare [62] to infrastructure [11] and manufacturing [31], underscores its pivotal role in modern analytics.
+These type of data are instrumental in identifying the patterns, dependencies, and anomalies indicative of significant shifts in system behaviors or the emergence of critical issues [12, 37].
+Traditionally, statistical methods have been the cornerstone for anomaly detection in time series data, and are renowned for their robust mathematical frameworks [60, 67, 49].
+However, the advent of Big Data, characterized by its significant volume, velocity, and variety, has revealed the limitations of these traditional methods, including high false positive rates and missed detections [3, 6].
+In response, deep learning has introduced a paradigm shift, offering enhanced adaptability and performance in anomaly detection tasks [27, 52, 26].
+Notably, the transformer architecture, with its self-attention mechanism, has emerged as a groundbreaking development, making significant strides not only in natural language processing but also in time series analysis [63, 59, 33].
+Despite the advantages offered by transformer models, their application to time series anomaly detection necessitates a more dynamic and adaptable framework for fine-tuning their structures and parameters for optimal performance [7].
+Our model, TransNAS-TSAD, combines the architectural strengths of transformers (e.g., TranAD) with the optimization strategy of a neural architecture search (NAS) to systematically explore a huge number of architectural configurations.
+The latter leverages the multi-objective optimization capabilities of the NSGA-II algorithm [22, 21], enabling an efficient exploration of complex search spaces [45, 47].
+This represents a substantial evolution of the TranAD’s static architecture, facilitating a balance between performance and computational efficiency.
+Adapting anomaly detection models to time series data is inherently complex, especially when addressing multivariate series that present high-dimensional challenges and necessitate sophisticated model architectures [68, 44].
+The selective attention capabilities of the transformer model offer a compelling solution, provided the architectures are meticulously tailored to the specific research questions and dataset characteristics [46].
+The convergence of NAS with multi-objective optimization algorithms like NSGA-II marks a strategic evolution in automated model design.
+This approach provides a methodical pathway to discovering optimal architectures that consider multiple, often conflicting, objectives, thereby enhancing the adaptability and effectiveness of anomaly detection models [36, 70].
+Our research is grounded on three fundamental principles designed to tackle the inherent challenges of time series anomaly detection.
+First, we emphasize the necessity of developing specialized models tailored to the unique temporal resolution and dimensional characteristics of time series data, ensuring they are finely attuned to the intricacies of temporal patterns.
+Second, we aim to maximize the utility of transformer architectures, exploiting their advanced capabilities for deep and nuanced interpretation of temporal data, which is critical for identifying subtle anomalies.
+Lastly, strategic optimization through Neural Architecture Search (NAS), particularly leveraging the NSGA-II algorithm, forms the cornerstone of our methodology.
+This approach allows us to systematically refine and perfect our models, ensuring they are not only highly effective but also optimized for the specific demands of the datasets they analyze.
+Together, these principles guide our pursuit of creating models that are both innovative and highly adapted to the complex domain of time series anomaly detection.
+Our contributions are delineated as follows:
+  • We introduce a novel approach that applies NAS with NSGA-II optimization to specifically fine-tune transformer architectures for time series anomaly detection, showcasing NAS’s adaptability in specialized domains.
+  • Our experimentation identifies key architectural configurations that elevate detection performance, highlighting the importance of architectural fine-tuning in enhancing model efficacy.
+  • We demonstrate substantial performance improvements across diverse datasets, proving our optimized models’ superiority over existing methods and illustrating the benefits of multi-objective optimization for anomaly detection.
+  • We strike an optimal balance between computational efficiency and analytical accuracy, presenting models that excel in real-time anomaly detection and are feasible for practical deployment, emphasizing efficiency in model design.
+  • The synergy between transformers’ structural strengths and NAS’s adaptive optimization creates uniquely effective models for time series anomaly detection, marking a significant advancement by integrating architectural innovation with strategic optimization.
 2 Related Work
-2.1 Anomaly Detection in Time Series Data
-Anomaly detection in time series data has been a focal point of research due to its critical applications across various domains such as finance, healthcare, and industrial automation.
-Traditional approaches have primarily relied on statistical methods, which assume specific data distributions or apply distance-based metrics (Leadbetter, (1991); Wang et al., (2011)).
-However, these methods often struggle with high-dimensional, complex time series data inherent in modern applications (Hodge & Austin, (2004); Patcha & Park, (2007)).
-With the advent of Big Data, researchers have sought more robust solutions capable of handling large volumes of high-velocity, high-variety data.
-Machine learning, and specifically deep learning, has emerged as a powerful approach to detect anomalies by learning representations of normal behavior and identifying deviations (Lee et al., (2022); Landauer et al., (2023); Al-amri et al., (2021)).
-2.2 Transformers in Time Series Analysis
-, The transformer architecture, introduced by Vaswani et al., (2017) has become a cornerstone for sequence modeling tasks, initially within the realm of natural language, processing and subsequently in time series analysis (Li et al., (2019); Tang & Matteson, (2021).
-Its self-attention mechanism allows it to capture temporal dependencies without the constraints of sequential data processing, distinguishing it from its predecessors like RNNs and LSTMs (Karita et al., (2019); Reza et al., (2022); Katrompas et al., (2022)).
-In the domain of time series anomaly detection, transformers have shown promise due to their ability to model complex patterns over time (Tuli et al., (2022); Liu et al., (2021); Wang et al., (2022)).
-However, the research is still nascent, with considerable scope for optimization and adaptation to various time series characteristics (Luo et al., (2021); Wen et al., (2022)).
-2.3 Neural Architecture Search (NAS)
-NAS has emerged as a paradigm to automate the design of neural network architectures.
-Early NAS approaches were computationally intensive, employing reinforcement learning or evolutionary algorithms to explore the design space (Liu et al., (2021); Chen et al., (2019); Chu et al., (2020)).
-The development of efficient NAS methods has been a recent trend, focused on reducing the computational resources required for the search process (Jin et al., (2019); He et al., (2020)).
-2.4 Multi-Objective Optimization in NAS
-Multi-objective optimization in NAS aims to find a trade-off between competing objectives, such as model performance and computational cost.
-NSGA-II, an evolutionary algorithm, has been widely used for this purpose due to its efficiency in finding a diverse set of Pareto optimal solutions (Wang et al., (2020); Lu et al., (2019); Ying et al., (2020)).
-The application of multi-objective optimization in NAS for time series anomaly detection is still an under-explored area, offering a fertile ground for research (Borchert et al., (2022); Lu et al., (2021)).
-2.5 Convergence of NAS and Transformers
-The intersection of NAS and transformer models represents the cutting edge of research in the quest for optimal architectures for specific tasks.
-In time series anomaly detection, this convergence promises architectures that are both performant and efficient, addressing the unique challenges of temporal data (Wen et al., (2022)).
-This work builds upon these foundations, aiming to address the identified gaps by leveraging the strengths of NAS and the transformer model to push the boundaries of anomaly detection in time series data.
+The quest to detect and understand anomalies within time series data spans various domains including, but not limited to, finance [10], healthcare [62], infrastructure [11], and manufacturing [31].
+This widespread interest underscores the fundamental role of anomaly detection in predictive maintenance, fraud detection, and system health monitoring, among other critical applications [12, 37].
+In finance, for example, anomaly detection can signal fraudulent transactions or market manipulations, while in healthcare, it might indicate abnormal patient conditions that require immediate attention [10].
+Initially, the field of time series anomaly detection heavily relied on statistical methods, grounded in robust mathematical frameworks, that have provided deep insights into data patterns and anomalies [49, 67].
+These methods, while effective for simpler datasets, often assume specific data distributions or employ distance-based metrics to identify outliers.
+However, their application to the high-dimensional, complex nature of multivariate time series data reveals significant limitations.
+The primary challenge lies not merely in capturing the correlations within the data, but in effectively managing the intricacies associated with multi-correlated variables.
+These complexities can lead to an increased rate of false positives and missed detections, as traditional methods struggle to adequately interpret the nuanced interactions and dependencies that characterize modern datasets [23, 56].
+The emergence of deep learning models, and their ability to learn high-level representations from massive amounts of data, have shown remarkable success in detecting anomalies in time series data [52].
+These models excel in identifying deviations from frequently occurring patterns, significantly reducing false positives and improving detection accuracy [35, 39, 5].
+A deeper exploration of machine learning models, including specific architectures and their successes across various applications, would provide valuable insights [36].
+Among the deep learning architectures that have gained prominence, the transformer architecture introduced by [63] has revolutionized how models process sequential data.
+Unlike its predecessors, such as RNNs and LSTMs, the transformer leverages a self-attention mechanism, enabling it to capture long-range dependencies and temporal patterns in data without being hindered by the sequential processing bottleneck [38, 58, 30, 57, 32].
+Its application in time series anomaly detection is particularly promising, offering a new perspective on modeling temporal anomalies.
+However, this area remains relatively unexplored, with significant potential for optimization and adaptation to the unique characteristics of time series data across different domains [60].
+The Neural Architecture Search (NAS) has emerged as a transformative approach in automating the design of neural network architectures [71, 47, 13].
+By systematically exploring a vast space of architectural configurations, NAS aims to identify models that achieve optimal performance for a given task.
+This automation is crucial in deep learning, where the design and tuning of models are both resource-intensive and require specialized knowledge [36, 18, 16].
+Recent advancements in NAS have focused on improving its efficiency, reducing the computational resources required for the search process, and making state-of-the-art model architectures accessible for a broader range of applications [43, 71].
+Incorporating multi-objective optimization into NAS, using algorithms like NSGA-II, represents a significant leap forward [21, 16].
+For time series anomaly detection, where models must process large volumes of data accurately and with computational efficiency, the ability to find a nuanced balance between these competing objectives is crucial [66, 42, 71].
+Despite the clear advantages, the application of multi-objective optimization in NAS for anomaly detection in time series is ripe for exploration and offers vast potential for groundbreaking research [36].
+Our work seeks to harness the strengths of Neural Architecture Search (NAS) and transformer models to directly address the unique challenges associated with temporal data, crafting models that are finely tuned for anomaly detection in time series.
+This integration aims to develop architectures that excel in both performance and efficiency, leveraging the latest advancements in the field [15].
+Given the nascent stage of combining NAS with transformer models for this purpose, our research not only contributes to the existing body of knowledge but also opens avenues for future exploration and innovation in anomaly detection methodologies [69].
+By situating our study within the context of ongoing research efforts and emerging trends, we aim to advance the state of the art in anomaly detection within time series data and contribute meaningfully to the broader discourse in this rapidly evolving field.
 3 Methodology
@@
 We commence with a clear definition of the problem, setting the stage for the subsequent methodological steps.
-This is followed by an in-depth discussion on data refinement techniques, ensuring that the data is appropriately pre-processed for our analysis.
-Next, we delve into the intricacies of our transformer architecture, highlighting its design and capabilities.
-A critical component of our approach is the neural architecture search (NAS) process, employing the NSGA-II algorithm for optimizing the transformer model, which we expound upon in detail.
-This is supplemented by an overview of the evolutionary process that iteratively refines our model architecture.
+This is followed by an in-depth discussion on data refinement techniques, ensuring that the data are appropriately pre-processed for our analysis.
+Next, we present the intricacies of our transformer architecture, highlighting its design and capabilities.
+A critical component of our approach is the neural architecture search (NAS), which employs the NSGA-II algorithm for optimizing the transformer model.
+That description is supplemented by an overview of the evolutionary process that iteratively refines our model architecture.
 We then transition to discussing advanced anomaly detection techniques, integrating adversarial elements into our model.
 The section culminates with a description of how we harness the full potential of TransNAS-TSAD for practical anomaly detection applications, illustrating the real-world implications of our research.
-Through this structured approach, we aim to provide a clear and logical flow of our methodology, ensuring that each step contributes coherently to achieving our research objectives.
 3.1 Problem Definition
 Our research pioneers the application of a multi-objective approach to anomaly detection in multivariate time series data, utilizing transformer models. The problem is defined as:
-Given a sequence of multivariate time series data $S=\{x_{1},x_{2},\ldots,x_{T}\}$, where each point $x_{t}$ is timestamped $t$ and resides in an $m$-dimensional space ($x_{t}\in\mathbb{R}^{m}$), our aims are:
+Given a sequence of multivariate time series data $S=\{x_{1},x_{2},\ldots,x_{T}\}$, where each point $x_{t}$ is timestamped $t$ and resides in an $m$-dimensional space ($x_{t}\in\mathbb{R}^{m}$), our aims are to perform:
   1. 1.
@@
     Anomaly Component Analysis: Our goal extends to identifying anomalous components within each data point of the series, generating a detailed prediction sequence $Y=\{y_{1},y_{2},\ldots,y_{T^{\prime}}\}$, where each $y_{t}$ pinpoints the specific anomalous dimensions at timestamp $t$.
-3.1.1 Multi-Objective Transformer Model Optimization
-At the heart of our methodology is the use of NSGA-II based neural architecture search (NAS) for optimizing the transformer architecture, focusing on:
-  • 1.
-    Maximizing Anomaly Detection Accuracy: We aim to enhance the F1 score for precise anomaly detection.
-  • 2.
+Multi-Objective Transformer Model Optimization
+At the heart of our methodology is the use of the NSGA-II based neural architecture search (NAS) for optimizing the transformer architecture, focusing on:
+  • Maximizing Anomaly Detection Accuracy: We aim to enhance the classification performance (F1 score) for precise anomaly detection.
   • Minimizing Architectural Complexity: Concurrently, reducing the model’s size and computational demands for efficiency in practical deployment.
 This dual-focused approach is a significant advancement in applying transformer models for anomaly detection.
-3.1.2 Data Processing and Model Training
-Prior to modeling, the time series data undergoes a preprocessing phase where normalization is applied to ensure uniformity of scale across all series components.
-The normalization process is formulated as follows:
-$x_{t}\leftarrow\frac{x_{t}-\min(S)}{\max(S)-\min(S)+\epsilon}$
-where $\min(S)$ and $\max(S)$ denote the component-wise minimum and maximum values observed in the training time series $S$, respectively.
+Data Processing and Model Training
+Prior to model training and classification, the time series data undergo a preprocessing phase where normalization is applied to ensure uniformity of scale across all series components.
+The normalization process is as follows:
+$x_{t}\leftarrow\frac{x_{t}-\min(S)}{\max(S)-\min(S)+\epsilon},$
+where $\min(S)$ and $\max(S)$ denote the component-wise minimum and maximum values observed in the training and testing time series data $S$, respectively.
 The term $\epsilon$ is a small constant introduced to prevent division by zero, ensuring numerical stability.
@@
 A context window of size $K$ is defined around each data point $x_{t}$, resulting in:
-$W_{t}=\{x_{t-K+1},\dots,x_{t}\}$
+$W_{t}=\{x_{t-K+1},\dots,x_{t}\},$
 where boundary cases are handled by replicating the first available data point to fill the window, maintaining a fixed size of $K$.
 This approach encapsulates the temporal dependencies inherent in time series data.
-3.1.3 Anomaly Scoring and Thresholding
+Anomaly Scoring and Thresholding
 Our model is designed to assign an anomaly score to each context window $W_{t}$, as opposed to direct binary labels.
@@
 A dynamic threshold $D$ is established based on the distribution of the scores computed on the training data, facilitating the distinction between normal and anomalous windows.
-The precise mechanism for determining $D$ will be further elaborated upon in the subsequent sections, focusing on its adaptive nature and how it relates to the detection accuracy.
+The precise mechanism for determining the dynamic threshold $D$ and how it relates to the detection accuracy are described in subsequent sections.
 3.2 TransNAS-TSAD Transformer Architecture
 In the TransNAS-TSAD framework, the transformer architecture is specially adapted for time series anomaly detection.
-This adaptation involves a unique configuration of the standard transformer model to align with the specific characteristics of time series data.
-Data Augmentation Mechanisms Within the Model: TransNAS-TSAD integrates specific data augmentation techniques within the model architecture.
-These techniques, which can be fine-tuned during training, include:
-  • 1.
+This involves a unique configuration of the standard transformer model to align with the specific characteristics of the time series data.
+Data Augmentation Mechanisms Within the Model:
+TransNAS-TSAD integrates specific data augmentation techniques within the model architecture. These techniques, which can be fine-tuned during training, include:
   • Gaussian Noise Augmentation (Optional): Adds Gaussian noise to the input sequence for variability and robustness.
-  • 2.
   • Time Warping & Time Masking Augmentation (Optional): Applies time warping or masking, enhancing the model’s generalization capabilities.
-Encoder Operations: The encoder in TransNAS-TSAD is designed with the following operations:
-  • 1.
+Encoder Operations
+The encoder in TransNAS-TSAD is designed with the following operations:
   • Linear Embedding (Optional): Transforms the input via linear mapping to a predefined dimensional space.
-  • 2.
   • Layer Normalization: Normalizes the sequence for stability in subsequent operations.
-  • 3.
   • Positional Encoding: Uses either sinusoidal or Fourier positional encoding, depending on the architecture optimized through NAS.
-  • 4.
   • Multi-Head Self Attention: A key feature in TransNAS-TSAD is the multi-head attention mechanism, which is aligned with the number of features in the data, inspired by the TranAD model.
 It employs the standard scaled-dot product attention mechanism:
-    $\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{\text{model}}}}\right)V$ (1)
+    $\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{\text{model}}}}\right)V.$ (1)
     The multi-head attention mechanism is formulated as:
-    $\text{MultiHeadAtt}(Q,K,V)=\text{Concat}(H_{1},H_{2},\ldots,H_{h})$ (2)
+    $\text{MultiHeadAtt}(Q,K,V)=\text{Concat}(H_{1},H_{2},\ldots,H_{h}),$ (2)
     where $H_{i}=\text{Attention}(Q_{i},K_{i},V_{i})$.
-  • 5.
   • Feedforward Neural Networks (FFNs) Post-Attention: Following the attention mechanism, the sequence is processed through a series of feedforward neural networks (FFNs).
@@
 This optimization, conducted using Optuna, is designed to balance anomaly detection performance with computational efficiency, ensuring that the FFN dimensions are optimally tuned for the specific demands of time series anomaly detection.
-Decoder Operations: The decoder mirrors the encoder’s complexity and includes additional transformations that integrate encoder memory, apply multi-head self-attention, and utilize feedforward neural networks, culminating in a final layer for accurate reconstruction.
+Decoder Operations
+The decoder mirrors the encoder’s complexity and includes additional transformations that integrate encoder memory, apply multi-head self-attention, and utilize feedforward neural networks, culminating in a final layer for accurate reconstruction.
 TransNAS-TSAD can utilize either a dual sequential decoder setup or an iterative approach, depending on the optimized configuration for maximum anomaly detection accuracy.
@@
 This step is essential in identifying an architecture that not only excels in anomaly detection accuracy but also aligns with the computational constraints of practical deployment scenarios.
-The rigorous optimization and selection of transformer architectures through NAS directly contribute to enhancing the performance and efficiency of the TransNAS-TSAD framework, thereby fulfilling our goal of developing a robust and adaptable anomaly detection system.
+The optimal selection of transformer architectures through NAS directly contributes to enhancing the performance and efficiency of the TransNAS-TSAD framework, thereby fulfilling our goal of developing a robust and adaptable anomaly detection system.
 3.3.1 Search Space Definition
-The search space outlined in Table 1 within TransNAS-TSAD represents a diverse and extensive array of potential model architectures and hyperparameters.
-This assortment is meticulously designed to tailor neural network configurations specifically for time-series anomaly detection:
-  • 1.
+The search space within TransNAS-TSAD and outlined in Table 1 represents a diverse and extensive array of potential model architectures and hyperparameters.
+This assortment is meticulously tailored toward neural network configurations designed specifically for time-series anomaly detection:
   • Architectural Parameters: These parameters encompass the number and types of layers in the encoder and decoder, optimized attention mechanisms for temporal data analysis, dimensions of feedforward networks, and a variety of positional encoding and normalization methods.
 Collectively, they enable the model to adapt its architecture for different characteristics of time-series data.
-  • 2.
   • Training Hyperparameters: This category includes a broad spectrum of hyperparameters, such as learning rates, batch sizes that are adaptable to varying computational resources, and dropout rates to ensure effective regularization.
@@
 The comprehensive nature of the search space in TransNAS-TSAD allows for the exploration and optimization of a wide range of models, ensuring that the most suitable architecture and hyperparameter settings are identified for effective anomaly detection in different time-series datasets.
-Parameter Type Parameter Name Search Space Description
-Training
-Learning Rate
-$1\times 10^{-5}$ to $1\times 10^{-1}$ (log scale)
-Rate at which the model learns.
-Dropout Rate
-0.1 to 0.5
-Regularization to prevent overfitting.
-Batch Size
-16 to 128 (step of 16)
-Number of samples per training step.
-Gaussian Noise
-$1\times 10^{-4}$ to $1\times 10^{-1}$ (log scale)
-Noise added for robustness.
-Time Warping
-True, False
-Augmentation technique for time series.
-Time Masking
-True, False
-Augmentation technique to mask intervals.
-Window Size
-10 to 30
-Length of the input sequence window for the model.
-Architectural
-Positional Encoding Type
-Sinusoidal, Fourier
-Encoding type for sequence position.
-Dimension Feedforward
-8 to 128 (log scale)
-Size of the feedforward network.
-Encoder Layers
-1 to 3
-Number of layers in the encoder.
-Decoder Layers
-1 to 3
-Number of layers in the decoder.
-Activation Function
-ReLU, Leaky ReLU, Sigmoid, Tanh
-Non-linearity after each layer.
-Attention Type
-Scaled Dot Product
-Type of attention mechanism.
-Number of Attention Heads
-Equal to feature dimension
-Parallel attention layers.
-Use Linear Embedding
-True, False
-Option to use a linear embedding layer.
-Layer Normalization
-Layer, Batch, Instance
-Type of normalization used.
-Self Conditioning
-True, False
-Conditioning strategy for the model.
-Number of FFN Layers
-1 to 3
-Layers in the feedforward network.
-Phase Type
-1phase, 2phase, Iterative
-Model’s reconstruction and refinement strategy.
 Table 1: Neural Architecture Search Space for TransNAS-TSAD
+Parameter Type Parameter Name Search Space Description Training Learning Rate $1\times 10^{-5}$ to $1\times 10^{-1}$ (log scale) Rate at which the model learns.
+Dropout Rate 0.1 to 0.5 Regularization to prevent overfitting.
+Batch Size 16 to 128 (step of 16) Number of samples per training step.
+Gaussian Noise $1\times 10^{-4}$ to $1\times 10^{-1}$ (log scale) Noise added for robustness.
+Time Warping True, False Augmentation technique for time series.
+Time Masking True, False Augmentation technique to mask intervals.
+Window Size 10 to 30 Length of the input sequence window for the model.
+Architectural Positional Encoding Type Sinusoidal, Fourier Encoding type for sequence position.
+Dimension Feedforward 8 to 128 (log scale) Size of the feedforward network.
+Encoder Layers 1 to 3 Number of layers in the encoder.
+Decoder Layers 1 to 3 Number of layers in the decoder.
+Activation Function ReLU, Leaky ReLU, Sigmoid, Tanh Non-linearity after each layer.
+Attention Type Scaled Dot Product Type of attention mechanism.
+Number of Attention Heads Equal to feature dimension Parallel attention layers.
+Use Linear Embedding True, False Option to use a linear embedding layer.
+Layer Normalization Layer, Batch, Instance Type of normalization used.
+Self Conditioning True, False Conditioning strategy for the model.
+Number of FFN Layers 1 to 3 Layers in the feedforward network.
+Phase Type 1phase, 2phase, Iterative Model’s reconstruction and refinement strategy.
 3.3.2 TransNAS-TSAD Evaluation Strategy
-Each trial within the TransNAS-TSAD NAS process entails a rigorous evaluation of models generated with varying architectural and training hyperparameters.
-This evaluation adopts a multi-objective approach, primarily focusing on two critical aspects: the F1 score and the number of model parameters.
+The evolutionary optimization strategy of TransNAS-TSAD entails a rigorous evaluation of models generated with varying architectural and training hyperparameters.
+The multi-objective evaluation focuses on two critical aspects: the F1 score and the number of model parameters.
 The F1 score serves as a key indicator of the model’s accuracy in anomaly detection, balancing precision and recall, while the number of parameters gauges the model’s architectural complexity and computational efficiency.
@@
 This dual-criteria assessment ensures the selection of models that are both effective in performance and practical in deployment, aligning with the overarching goal of achieving optimal anomaly detection with computational resourcefulness.
-3.3.3 Multi-Objective Optimization with NSGA-II
-TransNAS-TSAD employs the NSGA-II algorithm to balance the objectives of performance and computational efficiency. The optimization process is defined by the following equations:
-$\mathbf{F}(\mathbf{x})=[f_{1}(\mathbf{x}),f_{2}(\mathbf{x})]$ (3)
+Multi-Objective Optimization with NSGA-II
+TransNAS-TSAD employs the NSGA-II algorithm, a proven leader in balancing the objectives of performance and computational efficiency in multi-objective optimization.
+Our choice of NSGA-II is substantiated by recent benchmark studies, such as those presented in [45], which demonstrate its superiority in identifying optimal solutions across various search spaces.
+This evidence supports NSGA-II as an effective approach for optimizing the complex and competing demands of our model architecture.
+The optimization process in TransNAS-TSAD is governed by the following equations:
+$\mathbf{F}(\mathbf{x})=[f_{1}(\mathbf{x}),f_{2}(\mathbf{x})],$ (3)
 where $f_{1}(\mathbf{x})$ is the F1 score and $f_{2}(\mathbf{x})$ is a computational resource metric.
@@
 $d(\mathbf{x})=\sum_{i=1}^{k}\left(f_{i}(\mathbf{x}^{+})-f_{i}(\mathbf{x}^{-})\right)$ (5)
-3.3.4 Pareto Front Exploration
+Pareto Front Exploration and utilization
 The Pareto front represents the set of non-dominated solutions, providing an optimal trade-off between conflicting objectives:
-$\mathcal{P}=\{\mathbf{x}\in\mathbf{X}\mid\nexists\mathbf{y}\in\mathbf{X}:\mathbf{F}(\mathbf{y})\prec\mathbf{F}(\mathbf{x})\}$ (6)
+$\mathcal{P}=\{\mathbf{x}\in\mathbf{X}\mid\nexists\mathbf{y}\in\mathbf{X}:\mathbf{F}(\mathbf{y})\prec\mathbf{F}(\mathbf{x})\},$ (6)
 where $\prec$ indicates that $\mathbf{y}$ dominates $\mathbf{x}$.
-3.3.5 Utilization of the Pareto Front
 The Pareto front derived from the NAS process provides a practical framework for model selection:
-  • 1.
   • Interpretation and Analysis: The front is analyzed to discern architectural trade-offs, allowing stakeholders to identify models with the desired balance between accuracy and efficiency.
-  • 2.
   • Informed Selection: Decision-makers can select models that align with specific performance expectations and operational constraints by examining the Pareto front.
-  • 3.
   • Resource Allocation: Models on the Pareto front inform resource allocation, directing investments toward architectures with the most favorable trade-offs.
-Algorithm 1 provides a detailed procedure that encapsulates our approach within the Optuna framework using NSGA-II (Deb et al., (2002)) for the multi-objective optimization.
-This algorithm elucidates the sequential steps taken from initialization to final selection, highlighting the integration of the evolutionary search process with transformer-specific architecture and hyperparameter tuning.
-Input: Training dataset, validation dataset, predefined search space for transformer architectures and hyperparameters
-Output: Optimized transformer architecture and hyperparameters for time series anomaly detection
-1 Initialization: Set up the Optuna framework with the NSGA-II sampler.
+Algorithm 1 provides a detailed procedure that encapsulates our approach within the Optuna framework using NSGA-II [21] for the multi-objective optimization.
+This algorithm highlights, from initialization to final selection, the integration of the evolutionary search process with transformer-specific architecture and hyperparameter tuning.
+Algorithm 1 Multi-Objective NAS for Time Series Anomaly Detection with TransNAS-TSAD
+1:  Input: Training dataset, validation dataset, predefined search space for transformer architectures and hyperparameters
+2:  Output: Optimized transformer architecture and hyperparameters for time series anomaly detection
+3:  Initialization: Set up the Optuna framework with the NSGA-II sampler.
 Initialize the search with a population of transformer architectures from the predefined search space.
 Define evolutionary process parameters: population size, number of generations, mutation and crossover rates.
-Specify evaluation metrics: F1 score, computational resource utilization (model size, training time), precision, recall, and inference time.
-for each study in Optuna do
-2       for each trial in the study do
-3             Define the transformer model architectural parameters and training hyperparameters within the search space constraints.
-Construct and train the transformer model on the Training dataset.
-Evaluate the model on the Validation dataset using the specified metrics.
-Update the study with the trial’s objectives, reflecting the model’s performance and computational efficiency.
-if early stopping criteria are met based on validation performance or predefined resource constraints then
-4                   Terminate the trial prematurely.
-5             end if
-6
-7       end for
-8      Selection: Use Optuna’s NSGA-II sampler to select architectures for the next generation, focusing on the Pareto-optimal solutions.
-Crossover and Mutation: Apply Optuna’s genetic operators to evolve the study’s population, exploring new architectural variants.
-9 end for
-Final Selection: At the end of the NAS process, use Optuna to extract the best transformer architecture and hyperparameters from the Pareto front, considering both performance metrics and computational efficiency.
-Algorithm 1 Multi-Objective NAS for Time Series Anomaly Detection with TransNAS-TSAD
+4:  Specify evaluation metrics: F1 score, computational resource utilization (model size, training time), precision, recall, and inference time.
+5:  for each study in Optuna do
+6:     for each trial in the study do
+7:        Define the transformer model architectural parameters and training hyperparameters within the search space constraints.
+8:        Construct and train the transformer model on the Training dataset.
+9:        Evaluate the model on the Validation dataset using the specified metrics.
+10:        Update the study with the trial’s objectives, reflecting the model’s performance and computational efficiency.
+11:        if early stopping criteria are met based on validation performance or predefined resource constraints then
+12:           Terminate the trial prematurely.
+13:        end if
+14:     end for
+15:     Selection: Use Optuna’s NSGA-II sampler to select architectures for the next generation, focusing on the Pareto-optimal solutions.
+16:     Crossover and Mutation: Apply Optuna’s genetic operators to evolve the study’s population, exploring new architectural variants.
+17:  end for
+18:  Final Selection: At the end of the NAS process, use Optuna to extract the best transformer architecture and hyperparameters from the Pareto front, considering both performance metrics and computational efficiency.
 3.3.3 Evolutionary Process in TransNAS-TSAD
-The evolutionary process within TransNAS-TSAD serves a critical role in aligning our methodology with the overarching goal of refining anomaly detection capabilities.
+The evolutionary process within TransNAS-TSAD is critical in achieving our overarching goal of refining anomaly detection capabilities.
 Through iterative refinement and selection of model architectures, this process ensures that our models are not only accurate in identifying anomalies but also evolve to become more efficient and adaptable to various data characteristics.
 This continuous improvement and adaptation are key to achieving high-performance anomaly detection, which stands at the core of our research objectives.
-  • 1.
-    Generation Loop: New models are proposed, evaluated, and bred in each generation, with superior architectures refining the Pareto front.
-  • 2.
+  • Generation Loop: New models are generated, evaluated, and passed on to each generation, with superior architectures refining the Pareto front.
   • Early Stopping: The process incorporates early stopping for models that do not show promise, thus conserving computational resources.
-  • 3.
   • Final Architecture Selection: The optimal architecture is selected from the Pareto front, ensuring an effective balance between the objectives.
@@
   2. 2.
-    Achieving Pareto Front Efficiency: The key focus of NAS is to identify and refine a collection of architectures that constitute the Pareto front, representing a spectrum of models from high-performance, resource-intensive ones to balanced and resource-efficient alternatives.
+    Achieving Pareto Front Efficiency: The key focus of NAS is to identify and refine a collection of architectures that constitute the Pareto front, representing a spectrum of high-performance, resource-intensive models to more balanced and resource-efficient alternatives.
   3. 3.
     Iterative Evolutionary Process: Through evolutionary algorithms, these architectures undergo continuous adaptation and improvement, exploring innovative configurations and enhancing performance and efficiency with each generation.
-3.4.1 Tailoring Architectures to Specific Deployment Needs
+Tailoring Architectures to Specific Deployment Needs
 Post the offline NAS process, we are equipped with a diverse array of transformer architectures, each optimized for different operational contexts:
-  • 1.
   • For Resource-Rich Environments: High F1 score models, though computationally demanding, are suitable for scenarios where resource availability is not a constraint.
-  • 2.
   • In Resource-Limited Settings: The NAS process also yields architectures that are either balanced in terms of performance and resource usage or are specifically optimized for low-resource environments.
@@
 3.5 Advanced Anomaly Detection with Adversarial Elements in TransNAS-TSAD
-TransNAS-TSAD represents a significant advancement in the field of anomaly detection in time series data, strategically incorporating and extending adversarial learning paradigms.
-Drawing inspiration from the two-phase approach of the TranAD model, our framework elevates this concept by introducing a neural architecture search (NAS) driven methodology.
-This approach dynamically selects between the refined two-phase approach and our novel iterative self-adversarial approach at each trial.
-The integration of these advanced techniques addresses the complexities inherent in time series data, particularly in capturing subtle and sophisticated anomalies that traditional encoder-decoder architectures might overlook.
-Our development of the iterative self-adversarial approach, in particular, marks a notable enhancement over existing methods.
-By embedding elements of adversarial training within both the two-phase and iterative approaches, TransNAS-TSAD achieves a comprehensive and adaptable anomaly detection capability.
-This not only enhances the sensitivity and accuracy of the detection process but also ensures robustness and adaptability across diverse datasets.
-The synergy of these methodologies within TransNAS-TSAD aligns seamlessly with our research goal of developing a versatile and effective solution for time series anomaly detection, demonstrating our commitment to pushing the boundaries of current technologies.
-3.5.1 Two-Phase Adversarial Approach
-Drawing from adversarial learning paradigms, the two-phase approach in TransNAS-TSAD seeks to enhance the model’s sensitivity to anomalies through competitive reconstruction stages.
+TransNAS-TSAD represents a breakthrough in the field of time series anomaly detection by strategically enhancing adversarial learning paradigms.
+While it draws foundational inspiration from the TranAD model’s two-phase approach, our framework introduces: a neural architecture search (NAS) strategy that effectively incorporates a third, iterative self-adversarial phase.
+This tripartite approach enables dynamic selection among the conventional two-phase mechanism and our advanced iterative phase, optimizing detection capabilities for each specific trial.
+The inception of this third phase signifies a substantial evolution in anomaly detection techniques.
+By embedding adversarial elements within all three phases, TransNAS-TSAD not only retains the strengths of the traditional encoder-decoder models but also introduces a level of adaptability and precision previously unattainable.
+This method excels in identifying subtle and complex anomalies, leveraging the iterative adversarial training to refine detection with each iteration.
+3.5.1 Three-Phase Adversarial Approach
+Drawing from adversarial learning paradigms, the three-phase approach in TransNAS-TSAD seeks to enhance the model’s sensitivity to anomalies through competitive reconstruction stages.
 Phase 1 - Preliminary Input Reconstruction
-This phase, acting as the foundation, aims for a preliminary reconstruction of the input time-series window.
-The preliminary input reconstruction yields a focus score, defined by the deviation of the reconstructed output from the actual input:
-$L_{\text{focus}}=\|O_{\text{initial}}-W\|^{2}$
+This phase, acting as the foundation, aims for a preliminary reconstruction of the input time-series window, and yields a focus score.
+The latter is defined by the deviation of the reconstructed output from the actual input:
+$L_{\text{focus}}=\|O_{\text{initial}}-W\|^{2},$
 where $O_{\text{initial}}$ is the output from the first phase and $W$ is the input time-series window.
-The resultant “focus score,” derived from the deviations in this initial reconstruction, serves as an attention modulator for the subsequent phase.
+The resultant “focus score”, derived from the deviations in this initial reconstruction, serves as an attention modulator for the subsequent phase.
 Phase 2 - Adversarial, Focus-Driven Reconstruction
-Incorporating adversarial elements, this phase utilizes the focus scores from the phase 1 as a guiding beacon.
+Incorporating adversarial elements, this phase utilizes the focus scores from phase 1.
 In the adversarial phase, the second decoder aims to maximize the difference between its output and the input.
@@
 $L_{\text{adv1}}=\|O_{\text{adv1}}-W\|^{2}$
-$L_{\text{adv2}}=-\|O_{\text{adv2}}-W\|^{2}$
+$L_{\text{adv2}}=-\|O_{\text{adv2}}-W\|^{2},$
 where $O_{\text{adv1}}$ and $O_{\text{adv2}}$ are the outputs from the first and second decoders, respectively, during the adversarial phase.
-3.5.2 Iterative Self-Adversarial Approach
-Going beyond the structured two-phase reconstruction, TransNAS-TSAD introduces a dynamic iterative approach.
+Phase 3 -Iterative Self-Adversarial Approach
+Going beyond the structured two-phase reconstruction inspired by the TranAD work, TransNAS-TSAD introduces a dynamic iterative approach.
 This approach, embedded with self-adversarial mechanisms, continually refines its reconstructions.
-3.5.3 Iterative Refinement:
+Iterative Refinement:
 Starting with an initial reconstruction, the deviation between the current output and the input provides feedback for the next iteration:
@@
 The self-adversarial mechanism can be represented as:
-$L_{\text{self-adv}}=(L_{\text{iteration, prev}}-L_{\text{iteration, current}})^{2}$
+$L_{\text{self-adv}}=(L_{\text{iteration, prev}}-L_{\text{iteration, current}})^{2},$
 where $L_{\text{iteration, prev}}$ is the loss from the previous iteration and $L_{\text{iteration, current}}$ is the loss from the current iteration.
@@
 Mathematically, the convergence criterion can be defined as:
-$\Delta L=\left|L_{\text{iteration, current}}-L_{\text{iteration, prev}}\right|$
+$\Delta L=\left|L_{\text{iteration, current}}-L_{\text{iteration, prev}}\right|.$
 If $\Delta L<\epsilon$, where $\epsilon$ is a small positive value (e.g., $10^{-5}$), the iteration stops, suggesting that further refinement may not yield significant improvements.
@@
 The best score is derived from the iteration with the smallest reconstruction loss, ensuring that the selected representation most closely matches the input time-series data.
-3.5.4 Adaptable and Robust Anomaly Detection
+Adaptable and Robust Anomaly Detection
 What sets TransNAS-TSAD apart is its adaptability.
@@
 This section elucidates the methodologies employed to infer anomalies from time-series data, leveraging the trained transformer model and various augmentation strategies.
-3.6.1 The Anomaly Scoring Mechanism
+The Anomaly Scoring Mechanism
 Every incoming data point, denoted as $\hat{W}$, is assessed for anomalies by computing a score that quantitatively reflects its deviation from expected patterns.
-In this context, $\hat{W}$ represents the new window of the time series data under consideration, and $\hat{C}$ embodies any associated contextual information that could influence the anomaly detection process.
+Here, $\hat{W}$ represents the new window of time series data under consideration.
 The anomaly score, $s$, is calculated as:
-$s=\frac{1}{2}\|R_{1}-\hat{W}\|_{2}^{2}+\frac{1}{2}\|\hat{R}_{2}-\hat{W}\|_{2}^{2}$ (7)
+$s=\frac{1}{2}\|R_{1}-\hat{W}\|_{2}^{2}+\frac{1}{2}\|\hat{R}_{2}-\hat{W}\|_{2}^{2},$ (7)
 where $R_{1}$ and $\hat{R}_{2}$ are the reconstructed outputs from the transformer model, corresponding to different stages of the inference process.
-3.6.2 Dual Inference Pathways
+Dual Inference Pathways
 TransNAS-TSAD boasts of two distinct inference methodologies:
-  • 1.
   • Two-Phase Approach: A structured two-step inference process yielding reconstructions $(R_{1},\hat{R_{2}})$.
-  • 2.
   • Iterative Refinement: For scenarios demanding intricate attention, we engage in iterative inference, perfecting the anomaly scores with each iteration.
-Figure 2 illustrates the anomaly detection process on dimension 0 of the SMD test dataset using TransNAS-TSAD.
+Figure 2 illustrates the anomaly detection process on dimension 0 of the Server Machine dataset (SMD) test dataset using TransNAS-TSAD.
 The upper plot displays the true data points (in blue) and the predicted values (in red), while the lower plot indicates the computed anomaly scores and their corresponding labels, offering a visual interpretation of the model’s detection capabilities in identifying anomalies.
 Figure 2: Anomaly detection in TransNAS-TSAD on dimension 0 of the SMD test dataset, depicting true versus predicted values and identified anomalies
-3.6.3 Dynamic Anomaly Thresholding: The Sentinel
+Dynamic Anomaly Thresholding: The Sentinel
 Central to our thresholding strategy is the peaks over threshold (POT) method.
@@
 Evolutionary in nature, our modified POT (mPOT) approach ensures adaptability to the ever-changing landscape of time-series data:
-$mPOT(x)=POT(x)+\alpha\times\text{recent\_deviation}(x)$ (8)
+$mPOT(x)=POT(x)+\alpha\times\text{recent\_deviation}(x),$ (8)
 where $\alpha$ is a weight factor.
@@
 The moving average threshold at time $t$ is given by:
-$MAT(t)=\frac{1}{N}\sum_{i=t-N}^{t}s_{i}$ (9)
+$MAT(t)=\frac{1}{N}\sum_{i=t-N}^{t}s_{i},$ (9)
 where $(s_{i}$ is the anomaly score at time $i$ and $N$ represents the window size for the moving average.
@@
 For each data point at time $t$, the rolling mean $\mu(t)$ and rolling standard deviation $\sigma(t)$ are computed as:
-$\mu(t)=\frac{1}{W}\sum_{i=t-W}^{t}x_{i}\quad\text{and}\quad\sigma(t)=\sqrt{\frac{1}{W}\sum_{i=t-W}^{t}(x_{i}-\mu(t))^{2}}$ (10)
+$\mu(t)=\frac{1}{W}\sum_{i=t-W}^{t}x_{i}\quad\text{and}\quad\sigma(t)=\sqrt{\frac{1}{W}\sum_{i=t-W}^{t}(x_{i}-\mu(t))^{2}},$ (10)
 where $x_{i}$ represents the data point at time $i$ and $W$ is the rolling window size.
@@
 4 Experimental Setup for TransNAS-TSAD
-Our study rigorously evaluates the performance of TransNAS-TSAD against a suite of established benchmark models, including TranAD (Tuli et al., (2022)), LSTM-NDT (Hundman et al., (2018)), DAGMM (Zong et al., (2018)), OmniAnomaly (Su et al., (2019)), MSCRED (Zhang et al., (2019)), MAD-GAN (Li et al., (2019)), USAD (Audibert et al., (2020)), CAE-M (Zhang et al., (2021)), and GDN (Deng & Hooi, (2021)).
+We evaluate the performance of TransNAS-TSAD against a suite of established benchmark models, including TranAD [59], LSTM-NDT [24], DAGMM [72], OmniAnomaly [61], MSCRED [73], MAD-GAN [48], USAD [4], CAE-M [74]), and GDN [19].
 To ensure a fair comparison, we utilize the hyperparameter configurations as specified in the original publications of these models, relying on their publicly available implementations.
 The experimental infrastructure comprises a Google Colab Pro environment, leveraging a NVIDIA Tesla T4 GPU (16GB memory) and an Intel Xeon E5-2670 v3 CPU (8 cores, 51GB memory).
-The implementation is grounded in Python, utilizing robust libraries such as PyTorch for deep learning, Optuna for hyperparameter optimization, and pandas for data handling.
-These tools are instrumental in realizing the model implementation, executing the NSGA-II based NAS algorithm, and managing dataset operations.
-A pivotal element of our methodology is the neural architecture search (NAS), which automates the architectural design, navigating a comprehensive search space of architectural and training hyperparameters (detailed in Table 1).
-This search space is meticulously crafted to enable the NAS algorithm to adapt the model architecture to the unique characteristics of each dataset.
-Consistent with the practices in OmniAnomaly and TranAD, we apply the peaks over threshold (POT) method (Leadbetter, (1991)) across all datasets with a uniform coefficient of $10^{-4}$.
-The low quantile parameter, however, is fine-tuned for each specific dataset to maintain alignment with established benchmarks, thus ensuring an equitable basis for comparative analysis.
+The NSGA-II based NAS algorithm and data management operations are implemented in Python, utilizing robust libraries such as PyTorch for deep learning, Optuna for hyperparameter optimization, and pandas for data handling.
+A pivotal element of our methodology is the neural architecture search (NAS), which automates the architectural design by navigating a comprehensive search space of architectural and training hyperparameters, as shown in Table 1.
+This search space is crafted to enable the NAS algorithm to adapt the model architecture to the unique characteristics of each dataset.
+Consistent with the practices in OmniAnomaly and TranAD, we apply an enhanced version of the peaks over threshold (POT) method [49] across all datasets with a uniform coefficient of $10^{-4}$.
+This enhancement incorporates our previously described augmentative strategies, further fine-tuning the low quantile parameter for each specific dataset to align with established benchmarks and ensure equitable comparative analysis.
 4.1 Datasets
-Dataset Train Size Test Size Dimensions Sequences Anomalies (%)
-NAB 4,033 4,033 1 6 0.92
-UCR 1,600 5,900 1 4 1.88
-MBA 100,000 100,000 2 8 0.14
-SMAP 135,183 427,617 25 55 13.13
-MSL 58,317 73,729 55 3 10.72
-SWaT 496,800 449,919 51 1 11.98
-WADI 1,048,571 172,801 123 1 5.99
-SMD 708,405 708,420 38 4 4.16
 Table 2: Dataset Statistics
-Our study employs a comprehensive suite of datasets, each meticulously chosen to challenge and validate the robustness of time series anomaly detection methods across a variety of real-world scenarios.
-Table 2 summarizes the statistics of the datasets.
+Dataset Train Size Test Size Dimensions Sequences Anomalies (%) NAB 4,033 4,033 1 6 0.92 UCR 1,600 5,900 1 4 1.88 MBA 100,000 100,000 2 8 0.14 SMAP 135,183 427,617 25 55 13.13 MSL 58,317 73,729 55 3 10.72 SWaT 496,800 449,919 51 1 11.98 WADI 1,048,571 172,801 123 1 5.99 SMD 708,405 708,420 38 4 4.16
+Our study employs suite of datasets selected to challenge and validate the robustness of the time series anomaly detection methods across a variety of real-world scenarios.
 These datasets, renowned for their complexity and diversity, allow for an effective benchmarking of our TransNAS-TSAD approach against existing methodologies.
-  • 1.
-    Numenta Anomaly Benchmark (NAB) Dataset: OThis dataset encompasses a wide range of real-world data, including temperature sensor readings, cloud machine CPU utilization, service request latencies, and taxi demand data in New York City.
-It’s important to note some anomalies in labeling within this dataset, particularly in the nyc-taxi trace, which is often excluded from analyses (Ahmad et al., (2017); Nakamura et al., (2020)).
-  • 2.
+See Table 2 for summary statistics.
+  • Numenta Anomaly Benchmark (NAB) Dataset: This dataset encompasses a wide range of real-world data, including temperature sensor readings, cloud machine CPU utilization, service request latencies, and taxi demand data in New York City.
+It’s important to note some anomalies in labeling within this dataset are excluded from analyses, particularly in the NYC-taxi traces [1, 54].
   • HexagonML (UCR) Dataset: Featured in the KDD 2021 cup, this dataset comprises a diverse collection of univariate time series.
-In our research, we selectively utilized datasets that are derived from real-world sources, focusing specifically on the InternalBleeding and ECG datasets.
-This choice was driven by our need for natural data representations in our analysis.
-Consequently, we did not include any synthetic sequences that are also part of the HexagonML (UCR) Dataset (Dau et al., (2018)).”
-  • 3.
+In our research, we selectively use natural data representations derived from real-world sources, focusing specifically on the InternalBleeding and ECG datasets.
+Consequently, we did not include any synthetic sequences that are also part of the HexagonML (UCR) dataset [20].
   • MIT-BIH Supraventricular Arrhythmia Database (MBA) Dataset: This dataset includes electrocardiogram recordings from four patients, featuring anomalies such as supraventricular contractions or premature heartbeats.
-It’s a well-recognized dataset in data management studies (Moody & Mark, (2001); Boniol et al., (2020)).
-  • 4.
-    Soil Moisture Active Passive (SMAP) Dataset:: Provided by NASA, this dataset consists of soil samples and telemetry data from the Mars rover (Hundman et al., (2018)).
-  • 5.
-    Mars Science Laboratory (MSL) Dataset: Similar to the SMAP dataset, the MSL dataset includes sensor and actuator data from the Mars rover.(Hundman et al., (2018)) Due to the presence of many trivial sequences, only specific non-trivial sequences are typically analyzed such as (A4, C2 and T1) pointed by (Nakamura et al., (2020); Tuli et al., (2022)).
-  • 6.
+It’s a well-recognized dataset in data management studies [53, 9].
+  • Soil Moisture Active Passive (SMAP) Dataset: This NASA-provided dataset includes global measurements of soil moisture in the top 5 cm of Earth’s soil surface, collected approximately every three days by the SMAP satellite.
+It is designed to enhance our understanding of water, carbon, and energy cycles[24].
+  • Mars Science Laboratory (MSL) Dataset: Similar to the SMAP dataset, the MSL dataset includes sensor and actuator data from the Mars rover.[24] Due to the presence of many trivial sequences, only specific non-trivial sequences are typically analyzed such as (A4, C2 and T1) pointed by [54, 59].
   • Secure Water Treatment (SWaT) Dataset: This dataset is derived from a real-world water treatment plant, including data from 7 days of normal operations and 4 days under abnormal conditions.
-It features readings from various sensors and actuators (Mathur & Tippenhauer, (2016)).
-  • 7.
+It features readings from various sensors and actuators [51].
   • Water Distribution (WADI) Dataset: An expansion of the SWaT dataset, WADI includes a larger array of sensors and actuators.
-The dataset spans over a longer period, with 14 days of normal operation and 2 days under attack scenarios (Ahmed et al., (2017)).
-  • 8.
+The dataset spans a longer period, with 14 days of normal operation and 2 days under attack scenarios [2].
   • Server Machine Dataset (SMD): Covering five weeks of data, this dataset includes resource utilization traces from 28 machines in a compute cluster.
-Only specific non-trivial sequences are used for analysis (Su et al., (2019)).
-Each dataset is selected for its potential to challenge and validate the robustness of time series anomaly detection methods across various industry scenarios.
+Only specific non-trivial sequences are used for analysis [61].
 4.2 Evaluation Criteria
-In our study, aimed at advancing anomaly detection in time series data, we employ key metrics aligned with the objectives of the TransNAS-TSAD framework.
+We employ key metrics aligned with the objectives of the TransNAS-TSAD framework.
 These metrics are selected to optimize anomaly detection effectiveness, considering practical deployment aspects.
-4.2.1 F1 Score as Primary Performance Indicator
-The F1 score, the harmonic mean of precision and recall, is central to our evaluation strategy.
-It is particularly effective in addressing the challenge of imbalanced datasets common in anomaly detection, offering a balanced performance measure by accounting for the trade-off between false positives and false negatives.
+The F1 score, defined as the harmonic mean of precision and recall, is central to our evaluation strategy.
+It is a performance measure that balances the trade-off between false positives and false negatives to help address the challenge of imbalanced datasets common in anomaly detection.
 4.2.1 Introducing the Efficiency-Accuracy-Complexity Score (EACS)
-Model complexity is often indicated by parameter count.
-However, this metric can be misleading, especially for transformer-based models like TransNAS-TSAD.
-These models, despite their high parameter count, capture complex data patterns effectively, justifying their complexity.
-Therefore, we report parameter count and interpret it in the context of the model’s advanced capabilities.
-To comprehensively evaluate TransNAS-TSAD, we introduce the Efficiency-Accuracy-Complexity Score (EACS), a composite metric encompassing accuracy, training efficiency, and model complexity.
-It is defined as:
-$EACS=w_{\text{acc}}\times\left(\frac{\text{F1 Score}}{\text{Max F1 Score}}\right)+w_{\text{time}}\times\left(\frac{\text{Training Time}}{\text{Max Training Time}}\right)+w_{\text{params}}\times\left(\frac{\text{Parameter Count}}{\text{Max Parameter Count}}\right)$
-where $w_{acc}$, $w_{time}$, and $w_{params}$ are weights for the normalized F1 score, inverse training time, and inverse parameter count, respectively.
-Here, the weights are set as $w_{acc}=0.4$, $w_{time}=0.4$, and $w_{params}=0.2$, balancing accuracy, training efficiency, and model simplicity.
-This weighting scheme aligns with our research goals and practical deployment considerations, prioritizing accuracy and training efficiency while considering model complexity.
-Transformer models, known for their high parameter count, offer efficient processing capabilities through parallelization, making them suitable for scenarios requiring high accuracy where computational resources are not a constraint.
-4.3 Note on Additional Metrics
-Other metrics, such as ROC/AUC, are also valuable; however, our focus on the F1 score, parameter count, and EACS aligns with the specific goals of TransNAS-TSAD.
-This approach ensures a balanced evaluation of the models, considering both their effectiveness in anomaly detection and practical deployability.
+Model complexity is frequently measured by parameter count, a metric that, while useful,doesn’t fully capture the capabilities of transformer-based models like TransNASTSAD.
+Despite their high parameter count, these models are highly effective at capturing complex data patterns, a strength that justifies their complexity.
+Their advanced processing capabilities, facilitated by parallelization, make them adept at handling scenarios that demand high accuracy, even when computational resources are abundant.
+Therefore, when we report the parameter count of TransNAS-TSAD, it is within the broader context of its sophisticated functionality.
+Additionally, our evaluation emphasizes not just parameter count but also the F1 score and Efficiency- Accuracy-Complexity Score (EACS), ensuring our assessment considers both the model’s anomaly detection effectiveness and its deployability.
+While other metrics like ROC/AUC are valuable, our focused approach aligns with the specific objectives of TransNAS-TSAD, aiming for a balanced evaluation of its performance.
+[…]
+$EACS=w_{a}\times\left(\frac{\text{F1}}{\text{F1}_{\text{max}}}\right)+w_{t}\times\left(1-\frac{\text{T}_{\text{train}}}{\text{T}_{\text{max}}}\right)+w_{p}\times\left(1-\frac{\text{P}_{\text{count}}}{\text{P}_{\text{max}}}\right)$ (11)
+where: - $w_{a}$ is the weight for the accuracy component (F1 score), set to 0.4, - $w_{t}$ is the weight for the training time efficiency component, set to 0.4, - $w_{p}$ is the weight for the model complexity (parameter count) component, set to 0.2.
+These weights were chosen to reflect the importance of accuracy and training efficiency in practical deployments, where they are often prioritized over the complexity of the model.
+The F1 score (F1) represents the accuracy of the model, $\text{T}_{\text{train}}$ denotes the training time, and $\text{P}_{\text{count}}$ indicates the parameter count of the model.
 5 Results
-5.1 Evaluating TransNAS-TSAD: Precision, Recall, and F1 Score Analysis
-In our detailed evaluation of anomaly detection models, as captured in Table 3, the TransNAS-TSAD model emerges as a formidable contender, showcasing its prowess across diverse datasets including NAB, UCR, MBA, SMAP, MSL, SWaT, WADI, and SMD, with metrics of precision, recall, and F1 scores serving as the benchmarks.
-Standing out for its exceptional performance, TransNAS-TSAD notably surpasses other models in datasets such as NAB, UCR, MBA, and SMAP, with remarkable F1 scores (0.9411, 0.9910, 0.9861, and 0.9510 respectively), thereby demonstrating its adeptness in handling both univariate and multivariate time series.
-This is particularly evident when compared to models like OmniAnomaly and MSCRED, where TransNAS-TSAD’s ability to navigate complex data patterns becomes apparent.
-However, in the MSL dataset, the TranAD model, with its nuanced approach, shows a marginally superior performance with an F1 score of 0.9494 compared to TransNAS-TSAD’s 0.9482, possibly indicative of TranAD’s specific strengths in handling the dataset’s characteristics.
-Despite this, TransNAS-TSAD regains its lead in the SWaT dataset with an F1 score of 0.8314, highlighting its adaptability to various industrial contexts.
-Its exemplary performance continues in the challenging WADI and SMD datasets, where it achieves the highest F1 scores of 0.7735 and 0.9986, respectively, showcasing its unrivaled efficiency in detecting subtle anomalies in complex, high-dimensional datasets.
-This overarching success of TransNAS-TSAD, especially when contrasted with TranAD’s performance, underscores its comprehensive capability in anomaly detection.
-While models like LSTM-NDT and DAGMM show variable performance across different datasets, their inconsistency underlines TransNAS-TSAD’s steady effectiveness across various scenarios.
-Similarly, while OmniAnomaly and MSCRED perform well in certain contexts, they fall short of TransNAS-TSAD’s sophisticated temporal pattern handling.
-Other competitive models, including MAD-GAN and USAD, are eclipsed by TransNAS-TSAD’s more holistic approach, and even though CAE-M and GDN demonstrate strengths in specific scenarios, they generally do not match TransNAS-TSAD’s overall adaptability.
-The superiority of TransNAS-TSAD, especially in comparison to TranAD, invites further investigation through critical difference analysis, which would reinforce the statistical significance of the observed performance differences and provide a deeper understanding of TransNAS-TSAD’s robustness and reliability against a backdrop of diverse and challenging anomaly detection scenarios.
-Method NAB UCR MBA SMAP
-P R F1 P R F1 P R F1 P R F1
-LSTM-NDT 0.6400 0.6667 0.6531 0.5231 0.8294 0.5231 0.9207 0.9718 0.9456 0.8523 0.7326 0.7879
-DAGMM 0.7622 0.7292 0.7453 0.5337 0.9718 0.5337 0.9475 0.9900 0.9683 0.8069 0.9891 0.8888
-OmniAnomaly 0.8421 0.6667 0.7442 0.8346 0.9999 0.8346 0.8561 1.000 0.9225 0.813 0.9419 0.8728
-MSCRED 0.8522 0.6700 0.7502 0.5441 0.9718 0.5441 0.9272 1.0000 0.9623 0.8175 0.9216 0.8664
-MAD-GAN 0.8666 0.7012 0.7752 0.8538 0.9891 0.8538 0.9396 1.0000 0.9689 0.8157 0.9216 0.8654
-USAD 0.8421 0.6667 0.7442 0.8952 1.0000 0.8952 0.8953 0.9989 0.9443 0.748 0.9627 0.8419
-CAE-M 0.7918 0.8019 0.7968 0.6981 1.0000 0.6981 0.8442 0.9997 0.9154 0.8193 0.9567 0.8827
-GDN 0.8129 0.7872 0.7998 0.6894 0.9988 0.6894 0.8832 0.9892 0.9332 0.748 0.9891 0.8518
-TranAD 0.8889 0.9892 0.9364 0.9407 1.0000 0.9407 0.9569 1.0000 0.978 0.8043 0.9999 0.8915
-TransNAS-TSAD 0.8888 1.0000 0.9411 0.9823 1.0000 0.9910 0.9726 1.0000 0.9861 0.9066 1.0000 0.9510
-Method MSL SWaT WADI SMD
-P R F1 P R F1 P R F1 P R F1
-LSTM-NDT 0.6288 1.0000 0.7721 0.7778 0.5109 0.6167 0.0138 0.7823 0.0271 0.9736 0.844 0.9042
-DAGMM 0.7363 1.0000 0.8482 0.9933 0.6879 0.8128 0.076 0.9981 0.1412 0.9103 0.9914 0.9491
-OmniAnomaly 0.7848 0.9924 0.8765 0.9782 0.6957 0.8131 0.3158 0.6541 0.426 0.8881 0.9985 0.9401
-MSCRED 0.8912 0.9862 0.9363 0.9992 0.677 0.8072 0.2513 0.7319 0.3741 0.7276 0.9974 0.8414
-MAD-GAN 0.8516 0.993 0.9169 0.9593 0.6957 0.8065 0.2233 0.9124 0.3588 0.9991 0.844 0.915
-USAD 0.7949 0.9912 0.8822 0.9977 0.6879 0.8143 0.1873 0.8296 0.3056 0.906 0.9974 0.9495
-CAE-M 0.7751 1.0000 0.8733 0.9697 0.6957 0.8101 0.2782 0.7918 0.4117 0.9082 0.9671 0.9367
-GDN 0.9308 0.9892 0.9591 0.9697 0.6957 0.8101 0.2912 0.7931 0.426 0.717 0.9974 0.8342
-TranAD 0.9038 0.9999 0.9494 0.976 0.6997 0.8151 0.3529 0.8296 0.4951 0.9262 0.9974 0.9605
-TransNAS-TSAD 0.9567 1.000 0.94.82 0.9415 0.7624 0.8314 0.8626 0.7340 0.7735 0.9985 0.9988 0.9986
-Table 3: Benchmarking TransNAS-TSAD against Baseline Models: A summary of precision (P), recall (R), and F1 scores showcases the advanced anomaly detection capabilities of TransNAS-TSAD, derived via NSGA-II-based NAS, across diverse datasets.
-5.2 In-Depth Analysis of Training Efficiency and EACS in Anomaly Detection Models
-Furthering our investigation from Table 3, Table 4 provides a more granular look at the training efficiency and overall performance of anomaly detection models, with a particular emphasis on the Efficiency-Accuracy-Complexity Score (EACS).
-This table not only reiterates the superior performance of TransNAS-TSAD in precision and recall metrics, as observed in Table 3 but also introduces a critical analysis of training times and parameter counts across a range of datasets including NAB, UCR, MBA, SMAP, MSL, SWaT, WADI, and SMD.
-The training time, a pivotal yet often overlooked aspect, emerges as a decisive factor in evaluating the practicality of models.
-In operational settings, where rapid model deployment and iterative updating are essential, training time gains precedence over traditional metrics like parameter count.
-This is evidenced in the case of TransNAS-TSAD, which, across diverse datasets, demonstrates not only high F1 scores but also markedly reduced training times, leading to its high EACS.
-In the NAB dataset, for instance, TransNAS-TSAD achieves an F1 score of 94.11% with a training time of just 2.70 sec, resulting in an EACS of 0.9742.
-This trend of balancing accuracy with training efficiency is replicated in the UCR dataset, where it achieves an F1 score of 99.10% and an even lower training time of 2.24 sec, leading to an EACS of 0.9922.
-Diving deeper, in datasets like MBA and SMAP, TransNAS-TSAD’s EACS scores of 0.9932 and 0.9734, respectively, are not solely the product of its accuracy but are significantly bolstered by its training time efficiency.
-For example, in MBA, its training time of 2.27 is among the lowest, demonstrating the model’s capability to quickly adapt and learn from complex multivariate time series data.
-In comparison, models such as MSCRED, while exhibiting commendable F1 scores, are penalized in their EACS due to their disproportionately high training times (e.g., 774.99 sec for MBA), highlighting a crucial trade-off between model complexity and operational efficiency.
-In more challenging datasets like SWaT and WADI, the importance of training time becomes even more pronounced.
-TransNAS-TSAD’s ability to maintain high F1 scores while drastically cutting down on training times positions it as a highly effective tool for real-world applications.
-Its EACS of 0.9297 for SWaT and 0.6977 for WADI, while highlighting its proficiency in these industrial datasets, also reflect the balanced approach of the model, juxtaposing accuracy with swift adaptability.
-Furthermore, the analysis of models like GDN and TranAD provides additional context.
-While GDN shows strong performance in specific datasets like MSL with the highest F1 score (95.91%) and a respectable EACS of 0.9191, it falls behind TransNAS-TSAD in terms of training efficiency.
-Similarly, TranAD, despite its high F1 scores and relatively lower parameter counts, does not consistently match the training efficiency of TransNAS-TSAD, as seen in datasets like SMD and WADI.
-In conclusion, Table 4’s comprehensive analysis underscores the critical role of training time in the practical deployment of anomaly detection models.
-It establishes that models with lower training times, like TransNAS-TSAD, offer a significant advantage in real-world scenarios, effectively balancing accuracy, efficiency, and complexity.
-This detailed examination not only affirms the superiority of TransNAS-TSAD in standard performance metrics but also in its operational readiness, making it an exemplary model for efficient and effective anomaly detection in diverse and demanding environments.
-Method NAB UCR MBA SMAP
-F1 Training Time # Params EACS F1 Training Time # Params EACS F1 Training Time # Params EACS F1 Training Time # Params EACS
-LSTM-NDT 65.31 23.40 1710 0.8424 52.31 11.14 1765 0.7985 94.56 47.80 18641 0.9520 78.79 36.43 29146 0.8459
-DAGMM 74.53 64.50 1266 0.8467 53.37 27.50 1214 0.7919 96.83 92.36 2448 0.9394 88.80 26.60 7266 0.9050
-OmniAnomaly 74.42 88.70 13717 0.8250 83.46 38.21 13025 0.8863 92.25 197.66 25474 0.8649 87.28 66.77 16813 0.8230
-MSCRED 75.02 503.60 1237377 0.3001 54.41 559.60 128655 0.2176 96.23 774.99 2441778 0.3849 86.64 25.70 8237452 0.6982
-MAD-GAN 77.52 53.90 838 0.8671 85.38 31.50 822 0.9177 96.89 229.78 1877 0.8688 86.54 41.26 6718 0.8683
-USAD 74.42 65.40 1359 0.8455 89.52 54.23 1359 0.9172 94.43 191.26 1609 0.8789 84.19 49.22 7395 0.8439
-CAE-M 79.68 33.60 7229 0.8909 69.81 22.10 7365 0.8520 91.54 89.66 15411 0.9186 88.27 212.51 7229 0.5529
-GDN 79.98 131.60 574 0.8153 68.94 64.70 566 0.8286 93.32 203.45 1106 0.8682 85.18 104.50 2974 0.7440
-TranAD 93.64 3.37 615 0.9718 94.07 1.14 619 0.9745 97.80 5.11 1298 0.9885 89.15 5.62 62271 0.9445
-TransNAS-TSAD 94.11 2.70 285 0.9742 99.10 2.24 1690 0.9922 98.61 2.27 593 0.9932 94.60 3.19 39045 0.9734
-Method MSL SWaT WADI SMD
-F1 Training Time # Params EACS F1 Training Time # Params EACS F1 Training Time # Params EACS F1 Training Time # Params EACS
-LSTM-NDT 77.21 37.80 61856 0.8887 66.70 41.50 72854 0.7957 2.71 422.67 133712 0.5737 90.42 460.70 42853 0.9100
-DAGMM 84.82 25.66 18756 0.9258 81.28 29.66 16558 0.8748 14.12 246.32 39387 0.6392 94.91 337.90 10516 0.9420
-OmniAnomaly 87.65 29.69 39753 0.9348 81.31 40.20 36541 0.8568 42.60 361.60 39753 0.7471 94.01 311.20 22941 0.9412
-MSCRED 93.63 40.70 18476728 0.7535 80.72 236.80 14227264 0.3229 37.41 1884.25 75530 0.6435 84.14 349.77 14237356 0.6978
-MAD-GAN 91.69 32.64 17446 0.9497 80.65 34.70 15958 0.8638 35.88 321.20 35764 0.7228 91.50 424.60 9903 0.9188
-USAD 88.22 29.60 14859 0.9374 81.43 35.81 12502 0.8651 30.56 389.72 32859 0.6983 94.95 328.77 10609 0.9432
-CAE-M 87.33 774.60 204687 0.5471 81.01 71.22 194525 0.8010 41.17 7724.92 388905 0.3203 93.67 3606.90 114219 0.5731
-GDN 95.91 121.80 153541 0.9191 81.01 83.55 115442 0.7813 42.60 6047.12 287082 0.4245 83.42 1000.40 7255 0.8226
-TranAD 94.94 6.88 272181 0.9733 81.53 2.10 31336 0.9221 49.51 177.60 1378173 0.6316 96.05 56.70 135110 0.9760
-TransNAS-TSAD 94.82 5.44 236491 0.9739 83.14 1.60 13464 0.9297 77.35 225.60 1752919 0.6977 99.81 52.40 132050 0.9916
-Table 4: Comparative analysis of the Efficiency-Accuracy-Complexity Score (EACS) across various anomaly detection models, highlighting the performance of TransNAS-TSAD.
-The EACS is calculated by normalizing and weighting F1 scores (0.4), training times (0.4), and parameter counts (0.2), demonstrating the balanced efficiency and accuracy of TransNAS-TSAD against benchmark models across multiple datasets
-Arch and Training Params SMAP UCR MBA SWaT MSL SMD NAB WADI
-Learning Rate 0.0002128 0.0019997 0.003441 0.0000403 0.006192 0.0002273 0.006924 0.0000131
-Dropout Rate 0.2353 0.4474 0.1795 0.3836 0.3730 0.1554 0.4560 0.2237
-Dim Feedforward 101 124 41 42 121 86 10 16
-Batch Size 32 48 16 32 32 48 96 112
-Encoder Layers 2 1 3 2 2 3 1 2
-Decoder Layers 1 2 1 3 3 1 2 1
-Activation Func. sigmoid leaky_relu tanh tanh leaky_relu sigmoid relu sigmoid
-Time Warping False True True True True False False False
-Time Masking True True False True True False True False
-Gaussian Noise 0.000151 0.027956 0.007925 0.058019 0.000428 0.000628 0.000119 0.000194
-Linear Embedding False True False False True True True True
-Phase Type 2phase iterative iterative 2phase 2phase 2phase iterative 2Phase
-Self Conditioning False True False True True False True True
-Layer Norm False False False True True True False True
-Pos. Enc. Type sinusoidal fourier fourier sinusoidal fourier sinusoidal sinusoidal fourier
-FFN Layers 1 1 1 3 2 3 1 3
-Attn Heads 25 1 2 51 55 38 1 127
-Window Size 10 20 14 22 26 12 18 26
-Table 5: TransNAS-TSAD Hyperparameter Optimization Results: Detailed summary of the optimal training and architectural hyperparameters identified for each dataset, demonstrating the adaptive precision of TransNAS-TSAD in time series anomaly detection.
-5.2.1 Optimal Model Configurations Identified by TransNAS-TSAD
-Table 5 showcases the culmination of TransNAS-TSAD’s NSGA-II optimization process, highlighting the training and architectural hyperparameters of models that achieved the best balance between high F1 scores and low parameter counts.
-These configurations represent the pinnacle of the framework’s capability to identify models uniquely suited to each dataset’s anomaly detection needs.
-A key observation from the table is the diversity in architectural parameters across datasets, reflecting TransNAS-TSAD’s adeptness in tailoring models to specific data characteristics.
+Evaluating TransNAS-TSAD: Precision, Recall, and F1 Score Analysis
+The evaluation of anomaly detection models, as shown in Table LABEL:Table:3, measures TransNAS-TSAD’s performance across diverse datasets including NAB, UCR, MBA, SMAP, MSL, SWaT, WADI, and SMD, using precision, recall, and F1 scores as benchmarks.
+TransNAS-TSAD demonstrated high F1 scores in datasets such as NAB, UCR, MBA, and SMAP, benefiting from its advanced data processing techniques that adeptly handle complex patterns in both univariate and multivariate time series.
+This distinguishes it from models like OmniAnomaly and MSCRED.
+For the MSL dataset, the TranAD model slightly outperforms TransNAS-TSAD (0.9494 vs 0.9482), likely due to TranAD’s parameters being finely tuned to MSL’s unique characteristics.
+Despite conducting over 100 trials per dataset, this suggests that even more exhaustive optimization could potentially unlock further improvements for TransNAS-TSAD.
+Its robust performance in the SWaT dataset, with an F1 score of 0.8314, underscores its versatility in various industrial contexts.
+In the cases of WADI and SMD, where TransNAS-TSAD achieves F1 scores of 0.8400 and 0.9986 respectively, the significant improvements—such as a 40% increase over baseline in WADI—are attributable to its comprehensive search space and optimization strategies, enabling effective model tuning.
+This evaluation underscores the potential benefits of extending our optimization framework to achieve even greater model refinement.
+The variable performance of models like LSTM-NDT and DAGMM across different datasets points to their methodological strengths and constraints.
+For example, the LSTM-NDT’s reduced efficacy likely stems from the method’s nonparametric thresholding and resulting sensitivity to large differences in anomaly patterns; and the DAGMM struggles with longer sequences, in part, due to its reliance on a singular GRU model, which limits its capacity to accurately map extended temporal dynamics.
+Moreover, TransNAS-TSAD demonstrates enhanced performance over innovative yet less adaptable models such as MAD-GAN and USAD, and while CAE-M and GDN show potential in specific scenarios, they generally fall short of matching TransNAS-TSAD’s applicability and performance across the datasets reviewed, highlighting the value of integrating adversarial training and iterative optimization into the TransNAS-TSAD’s framework.
+, derived via NSGA-II-based NAS, across diverse datasets.,label=tab:Benchmarking,nofloat=table Method NAB UCR MBA SMAP P R F1 P R F1 P R F1 P R F1 LSTM-NDT 0.6400 0.6667 0.6531 0.5231 0.8294 0.5231 0.9207 0.9718 0.9456 0.8523 0.7326 0.7879 DAGMM 0.7622 0.7292 0.7453 0.5337 0.9718 0.5337 0.9475 0.9900 0.9683 0.8069 0.9891 0.8888 OmniAnomaly 0.8421 0.6667 0.7442 0.8346 0.9999 0.8346 0.8561 1.000 0.9225 0.813 0.9419 0.8728 MSCRED 0.8522 0.6700 0.7502 0.5441 0.9718 0.5441 0.9272 1.0000 0.9623 0.8175 0.9216 0.8664 MAD-GAN 0.8666 0.7012 0.7752 0.8538 0.9891 0.8538 0.9396 1.0000 0.9689 0.8157 0.9216 0.8654 USAD 0.8421 0.6667 0.7442 0.8952 1.0000 0.8952 0.8953 0.9989 0.9443 0.748 0.9627 0.8419 CAE-M 0.7918 0.8019 0.7968 0.6981 1.0000 0.6981 0.8442 0.9997 0.9154 0.8193 0.9567 0.8827 GDN 0.8129 0.7872 0.7998 0.6894 0.9988 0.6894 0.8832 0.9892 0.9332 0.748 0.9891 0.8518 TranAD 0.8889 0.9892 0.9364 0.9407 1.0000 0.9407 0.9569 1.0000 0.978 0.8043 0.9999 0.8915 TransNAS-TSAD 0.8888 1.0000 0.9411 0.9823 1.0000 0.9910 0.9726 1.0000 0.9861 0.9066 1.0000 0.9510 Method MSL SWaT WADI SMD P R F1 P R F1 P R F1 P R F1 LSTM-NDT 0.6288 1.0000 0.7721 0.7778 0.5109 0.6167 0.0138 0.7823 0.0271 0.9736 0.844 0.9042 DAGMM 0.7363 1.0000 0.8482 0.9933 0.6879 0.8128 0.076 0.9981 0.1412 0.9103 0.9914 0.9491 OmniAnomaly 0.7848 0.9924 0.8765 0.9782 0.6957 0.8131 0.3158 0.6541 0.426 0.8881 0.9985 0.9401 MSCRED 0.8912 0.9862 0.9363 0.9992 0.677 0.8072 0.2513 0.7319 0.3741 0.7276 0.9974 0.8414 MAD-GAN 0.8516 0.993 0.9169 0.9593 0.6957 0.8065 0.2233 0.9124 0.3588 0.9991 0.844 0.915 USAD 0.7949 0.9912 0.8822 0.9977 0.6879 0.8143 0.1873 0.8296 0.3056 0.906 0.9974 0.9495 CAE-M 0.7751 1.0000 0.8733 0.9697 0.6957 0.8101 0.2782 0.7918 0.4117 0.9082 0.9671 0.9367 GDN 0.9308 0.9892 0.9591 0.9697 0.6957 0.8101 0.2912 0.7931 0.426 0.717 0.9974 0.8342 TranAD 0.9038 0.9999 0.9494 0.976 0.6997 0.8151 0.3529 0.8296 0.4951 0.9262 0.9974 0.9605 TransNAS-TSAD 0.9567 1.000 0.94.82 0.9415 0.7624 0.8314 0.8508 0.8295 0.8400 0.9985 0.9988 0.9986
+In-Depth Analysis of Training Efficiency and EACS in Anomaly Detection Models
+Table LABEL:Tab:_EACS presents a comprehensive analysis that compares the Efficiency-Accuracy-Complexity Score (EACS) among various anomaly detection models, including TransNAS-TSAD.
+It highlights how training time and model size significantly influence the application of anomaly detection models in operational environments that demand rapid deployment and updates, a factor underscored by several studies [29, 55, 17].
+In this analysis, TransNAS-TSAD emerges as a standout performer, achieving high EACS values across multiple datasets, which indicates its effective balance of model performance and operational efficiency.
+Specifically, TransNAS-TSAD delivers high F1 scores of 94.11% and 99.10% on the NAB and UCR datasets, respectively, while maintaining brief training periods of 2.70 and 2.24 seconds.
+This translates into EACS values of 0.9742 and 0.9922, showcasing TransNAS-TSAD’s rapid adaptability and ability to learn from complex data.
+The trend (i.e., high accuracies combined with short training times) is consistent across other datasets, including MBA and SMAP, where TransNAS-TSAD records EACS scores of 0.9932 and 0.9734, respectively.
+This performance sets a stark contrast to other models, such as MSCRED, which achieves commendable F1 scores, but experiences a notable decrease in EACS due to prolonged training times (e.g., 774.99 seconds for MBA).
+In this light, TransNAS-TSAD distinguishes itself by adeptly optimizing both model complexity and training efficiency, showcasing its ability to balance high-performance metrics with operational practicality.
+The latter is particularly evident in datasets with challenging environments like SWaT and WADI, where TransNAS-TSAD not only sustains high F1 scores but also significantly reduces training times, achieving EACS scores of 0.9297 and 0.7883, respectively.
+While models like GDN and TranAD perform well in specific datasets, they do not consistently match TransNAS-TSAD in training efficiency across the board.
+This evaluation emphasizes the importance of considering training time and model size when deploying practical anomaly detection models.
+.
+The EACS is calculated by normalizing and weighting F1 scores (0.4), training times (0.4), and parameter counts (0.2), demonstrating the balanced efficiency and accuracy of TransNAS-TSAD against benchmark models across multiple datasets,label=Tab: EACS,nofloat=table Method NAB UCR MBA SMAP F1 % Training Time (sec) # Params EACS F1 % Training Time (sec) # Params EACS F1 % Training Time (sec) # Params EACS F1 % Training Time (sec) # Params EACS LSTM-NDT 65.31 23.40 1710 0.8424 52.31 11.14 1765 0.7985 94.56 47.80 18641 0.9520 78.79 36.43 29146 0.8459 DAGMM 74.53 64.50 1266 0.8467 53.37 27.50 1214 0.7919 96.83 92.36 2448 0.9394 88.80 26.60 7266 0.9050 OmniAnomaly 74.42 88.70 13717 0.8250 83.46 38.21 13025 0.8863 92.25 197.66 25474 0.8649 87.28 66.77 16813 0.8230 MSCRED 75.02 503.60 1237377 0.3001 54.41 559.60 128655 0.2176 96.23 774.99 2441778 0.3849 86.64 25.70 8237452 0.6982 MAD-GAN 77.52 53.90 838 0.8671 85.38 31.50 822 0.9177 96.89 229.78 1877 0.8688 86.54 41.26 6718 0.8683 USAD 74.42 65.40 1359 0.8455 89.52 54.23 1359 0.9172 94.43 191.26 1609 0.8789 84.19 49.22 7395 0.8439 CAE-M 79.68 33.60 7229 0.8909 69.81 22.10 7365 0.8520 91.54 89.66 15411 0.9186 88.27 212.51 7229 0.5529 GDN 79.98 131.60 574 0.8153 68.94 64.70 566 0.8286 93.32 203.45 1106 0.8682 85.18 104.50 2974 0.7440 TranAD 93.64 3.37 615 0.9718 94.07 1.14 619 0.9745 97.80 5.11 1298 0.9885 89.15 5.62 62271 0.9445 TransNAS-TSAD 94.11 2.70 285 0.9742 99.10 2.24 1690 0.9922 98.61 2.27 593 0.9932 94.60 3.19 39045 0.9734 Method MSL SWaT WADI SMD F1 % Training Time (sec) # Params EACS F1 Training Time (sec) # Params EACS F1 % Training Time (sec) # Params EACS F1 % Training Time (sec) # Params EACS LSTM-NDT 77.21 37.80 61856 0.8887 66.70 41.50 72854 0.7957 2.71 422.67 133712 0.5753 90.42 460.70 42853 0.9100 DAGMM 84.82 25.66 18756 0.9258 81.28 29.66 16558 0.8748 14.12 246.32 39387 0.6498 94.91 337.90 10516 0.9420 OmniAnomaly 87.65 29.69 39753 0.9348 81.31 40.20 36541 0.8568 42.60 361.60 39753 0.7794 94.01 311.20 22941 0.9412 MSCRED 93.63 40.70 18476728 0.7535 80.72 236.80 14227264 0.3229 37.41 1884.25 75530 0.6716 84.14 349.77 14237356 0.6978 MAD-GAN 91.69 32.64 17446 0.9497 80.65 34.70 15958 0.8638 35.88 321.20 35764 0.7500 91.50 424.60 9903 0.9188 USAD 88.22 29.60 14859 0.9374 81.43 35.81 12502 0.8651 30.56 389.72 32859 0.7214 94.95 328.77 10609 0.9432 CAE-M 87.33 774.60 204687 0.5471 81.01 71.22 194525 0.8010 41.17 7724.92 388905 0.3503 93.67 3606.90 114219 0.5731 GDN 95.91 121.80 153541 0.9191 81.01 83.55 115442 0.7813 42.60 6047.12 287082 0.4559 83.42 1000.40 7255 0.8226 TranAD 94.94 6.88 272181 0.9733 81.53 2.10 31336 0.9221 49.51 177.60 1378173 0.6645 96.05 56.70 135110 0.9760 TransNAS-TSAD 94.82 5.44 236491 0.9739 83.14 1.60 13464 0.9297 84.35 227.50 1701515 0.7883 99.81 52.40 132050 0.9916
+in time series anomaly detection.,label=tab:Best model params,nofloat=table Arch and Training Params SMAP UCR MBA SWaT MSL SMD NAB WADI Learning Rate 0.0002128 0.0019997 0.003441 0.0000403 0.006192 0.0002273 0.006924 0.0015406 Dropout Rate 0.2353 0.4474 0.1795 0.3836 0.3730 0.1554 0.4560 0.2886 Dim Feedforward 101 124 41 42 121 86 10 24 Batch Size 32 48 16 32 32 48 96 128 Encoder Layers 2 1 3 2 2 3 1 2 Decoder Layers 1 2 1 3 3 1 2 1 Activation Func.
+sigmoid leaky_relu tanh tanh leaky_relu sigmoid relu leaky_relu Time Warping False True True True True False False True Time Masking True True False True True False True False Gaussian Noise 0.000151 0.027956 0.007925 0.058019 0.000428 0.000628 0.000119 0.001042 Linear Embedding False True False False True True True True Phase Type 2phase iterative iterative 2phase 2phase 2phase iterative 2Phase Self Conditioning False True False True True False True False Layer Norm False False False True True True False False Pos.
+Enc.
+Type sinusoidal fourier fourier sinusoidal fourier sinusoidal sinusoidal sinusoidal FFN Layers 1 1 1 3 2 3 1 1 Attn Heads 25 1 2 51 55 38 1 127 Window Size 10 20 14 22 26 12 18 26
+5.1 Optimal Model Configurations Identified by TransNAS-TSAD
+Table LABEL:tab:Best_model_params showcases the culmination of TransNAS-TSAD’s NSGA-II optimization process, highlighting the training and architectural hyperparameters of models that achieved the best balance between high F1 scores and low parameter counts.
+These configurations represent models uniquely suited to each dataset’s anomaly detection needs.
+A key observation is the diversity in architectural parameters across datasets, reflecting TransNAS-TSAD’s adeptness in tailoring models to specific data characteristics.
 For example, in the SMAP dataset, the model employs a relatively simple architecture with fewer encoder and decoder layers, which is effective in handling the environmental time series data of this dataset.
@@
 For instance, the sinusoidal positional encoding in SMAP and NAB caters to their unique temporal patterns, while the Fourier positional encoding in UCR and WADI aligns with the datasets’ complex spectral characteristics.
-These results from TransNAS-TSAD’s NAS process validate the framework’s ability to not only fine-tune hyperparameters for operational efficiency but also to intricately adapt its architectural design to the nuanced demands of various time series anomaly detection scenarios, ensuring optimal detection performance across a broad spectrum of datasets.
+These results from TransNAS-TSAD’s NAS process validate the framework’s ability to not only fine-tune hyperparameters for operational efficiency, but also to intricately adapt its architectural design to the nuanced demands of various time series anomaly detection scenarios, ensuring optimal detection performance across a broad spectrum of datasets.
 Figure 3: Analysis of hyperparameter importance for F1 score optimization in TransNAS-TSAD across four datasets (NAB, MBA, SMAP, and WADI).
@@
 The MBA dataset’s analysis underscores the importance of activation_function as 0.11, num_feedforward_layers as 0.07, and dropout_rate as 0.07, indicating the need for a complex model architecture to effectively learn from multivariate ECG time series.
-In the SMAP dataset, gaussian_noise with importance 0.53 is paramount, highlighting the model’s need to handle noisy environmental data effectively.
+In the SMAP dataset, gaussian_noise with importance 0.53 highlights the model’s ability to handle noisy environmental data effectively.
 Additionally, batch_size with importance 0.07 influences the model’s performance, reflecting the impact of batch processing on training.
@@
 This pattern reflects the impact of data quality and structure on the model’s learning efficacy.
-The conducted hyperparameter analysis illuminates the tailored performance of TransNAS-TSAD across varied datasets, with each dataset’s specifics dictating the importance of particular hyperparameters.
+This hyperparameter analysis illuminates the tailored performance of TransNAS-TSAD across varied datasets, with each dataset’s specifics dictating the importance of particular hyperparameters.
 This not only affirms the model’s adaptability but also its capacity for ongoing enhancement, ensuring it remains adept at confronting the evolving complexities of time series data.
@@
 5.3 Pareto Front Analysis for Model Optimization
-Building upon the insights from the hyperparameter importance analysis, the Pareto front optimization results for the TransNAS-TSAD model, as depicted in Figure 4, further elucidate the model’s performance.
+Building upon the insights from the hyperparameter importance analysis, the Pareto front optimization results for the TransNAS-TSAD model are depicted in Figure 4.
 The Pareto front plots for the NAB, MBA, SMAP, and WADI datasets underscore the delicate balance between model complexity, as measured by the number of parameters, and the model’s effectiveness, as quantified by the F1 score.
@@
 A significant number of these configurations exhibit a high parameter count without a corresponding increase in the F1 score, suggesting a potential plateau in performance gains relative to complexity.
-This observation prompts a critical evaluation of model parsimony, encouraging the selection of simpler models that maintain performance while mitigating the risk of overfitting.
+This observation prompts a critical evaluation of model parsimony, encouraging the selection of simpler models that maintain performance while mitigating the risk of over-fitting.
 In the MBA dataset, the dispersion of trials across the Pareto front reflects a comprehensive exploration of the architectural space.
@@
 These Pareto fronts provide a visual and quantitative tool for identifying optimal model configurations that achieve a balance between performance and complexity.
-They serve as a decision-making aid for selecting models that not only excel in accuracy but are also efficient and scalable, aligning with the practical demands of anomaly detection in diverse environments.
-The findings from the Pareto front analysis are integral to the ongoing development of the TransNAS-TSAD model within the broader scope of our research.
-They contribute to the understanding of how different model architectures perform across various datasets, offering a pathway to enhancing the model’s adaptability and ensuring its continued efficacy in the face of evolving data challenges.
+They serve as a decision-making aid for selecting models that align with the practical demands of anomaly detection in diverse environments.
+The findings from the Pareto analysis are integral to the ongoing development of the TransNAS-TSAD model within the broader scope of our research.
+They contribute to understanding how different model architectures perform across various datasets, and offer a pathway to enhancing the model’s adaptability and ensuring its continued efficacy in the face of evolving data challenges.
 6 Discussion: Addressing Challenges and Exploring Future Directions in TransNAS-TSAD
@@
 This confluence of technologies marks a substantial leap forward, particularly in the detection of subtle and complex anomalies that have typically eluded traditional methods.
-A pivotal challenge addressed by TransNAS-TSAD is the balance of detection accuracy against computational efficiency.
-The NSGA-II algorithm plays a crucial role in this regard, effectively managing the equilibrium to ensure optimal model performance, without incurring excessive computational demands.
+A pivotal challenge addressed by TransNAS-TSAD is the balance between detection accuracy and computational efficiency.
+The NSGA-II algorithm plays a critical role in optimizing model performance, without incurring excessive computational demands.
 This aspect is particularly crucial in real-world applications where resources are finite and efficiency is paramount.
@@
 Additionally, the development and implementation of dynamic thresholding strategies, such as the Moving Average Thresholding (MAT) and the incorporation of rolling statistics, offer exciting prospects.
-These augmentative strategies could pave the way for even more nuanced detection capabilities within the TransNAS-TSAD framework.
 Furthermore, the exploration of hybrid systems that synergize machine learning with simulation approaches, along with advancements in neuro-symbolic systems, could substantially enhance the model’s adaptability and effectiveness across various scenarios.
@@
 7 Conclusion
-TransNAS-TSAD emerges as a groundbreaking framework in the realm of time series anomaly detection, effectively integrating transformer architecture with neural architecture search and NSGA-II optimization.
-The model’s superior performance, evidenced across diverse datasets, demonstrates its robustness in accurately detecting anomalies in both univariate and multivariate time series data.
-The strategic application of NSGA-II within TransNAS-TSAD skillfully addresses the challenge of balancing accuracy with computational efficiency, a critical consideration in practical applications.
-Moreover, the incorporation of advanced adversarial learning paradigms contributes to its ability to detect nuanced anomalies, a notable advancement over traditional methods.
-TransNAS-TSAD’s development signifies a notable contribution to the field of anomaly detection, enhancing the capabilities and efficiency of current methodologies.
-As the landscape of data analysis continues to evolve, the principles and approaches embodied in TransNAS-TSAD will undoubtedly inspire further advancements in machine learning, paving the way for innovative applications in various sectors.
-The framework not only sets a new benchmark in anomaly detection but also opens up exciting prospects for future research, particularly in enhancing real-time processing and integrating human-centric approaches to machine learning.
+TransNAS-TSAD signifies a notable contribution in the realm of time series anomaly detection by merging transformer architecture with neural architecture search and NSGA-II optimization, achieving superior performance across diverse univariate and multivariate datasets.
+Its robustness in accurately detecting anomalies highlights the effectiveness of this integration, effectively addressing the dual challenge of maintaining accuracy while ensuring computational efficiency—key for practical applications.
+This framework is further distinguished by its incorporation of advanced adversarial learning paradigms, enabling the detection of nuanced anomalies and marking a significant step forward from traditional methods.
+As the landscape of data analysis evolves, TransNAS-TSAD not only sets new performance benchmarks in anomaly detection but also opens up exciting avenues for future research, particularly in enhancing real-time processing and integrating human-centric approaches to machine learning.
+The principles and approaches embodied in TransNAS-TSAD are paving the way for innovative applications across various sectors, shaping the future of machine learning with its groundbreaking methodologies.
 Acknowledgments
@@
 We would like to extend our gratitude to the providers of the data sets used in this work, whose contributions were invaluable to our research.
-The authors would like to thank the US Geological Survey (USGS) for offering the domain expertise that was crucial to identify the peak anomaly types that are of practical importance.